### v4

# Quantum-Classical Hybrid Quantized Neural Network

Abstract
In this work, we introduce a novel Quadratic Binary Optimization (QBO) framework for training a quantized neural network. The framework enables the use of arbitrary activation and loss functions through spline interpolation, while Forward Interval Propagation addresses the nonlinearities and the multi-layered, composite structure of neural networks via discretizing activation functions into linear subintervals. This preserves the universal approximation properties of neural networks while allowing complex nonlinear functions accessible to quantum solvers, broadening their applicability in artificial intelligence. Theoretically, we derive an upper bound on the approximation error and the number of Ising spins required by deriving the sample complexity of the empirical risk minimization problem from an optimization perspective. A key challenge in solving the associated large-scale Quadratic Constrained Binary Optimization (QCBO) model is the presence of numerous constraints. When employing the penalty method to handle these constraints, tuning a large number of penalty coefficients becomes a critical hyperparameter optimization problem, increasing computational complexity and potentially affecting solution quality. To overcome this, we adopt the Quantum Conditional Gradient Descent (QCGD) algorithm, which solves QCBO directly on quantum hardware. We establish the convergence of QCGD under a quantum oracle subject to randomness, bounded variance, and limited coefficient precision, and further provide an upper bound on the Time-To-Solution. To enhance scalability, we further incorporate a decomposed copositive optimization scheme that replaces the monolithic lifted model with sample-wise subproblems. This decomposition substantially reduces the quantum resource requirements and enables efficient low-bit neural network training. We further propose the usage of QCGD and Quantum Progressive Hedging (QPH) algorithm to efficiently solve the decomposed problem. Experimental results using a coherent Ising machine achieve $94.95\%$ accuracy on the Fashion MNIST classification task with only 1.1-bit precision. Compared to classical quantization-aware training, post-training quantization methods, and a full precision model, our approach delivers higher accuracy, faster training, demonstrating its overall efficiency. Additionally, evaluations show substantial improvements in memory footprint and inference latency with negligible accuracy loss. We further validate the robustness of the spline interpolation method and analyze the effect of precision levels on QCGD convergence, confirming the reliability of the proposed framework.
Introduction
In recent years, the rapid progress of deep learning has catalyzed significant advances in a wide array of domains, from natural language processing [1] to computer vision [2], allowing intelligent systems to permeate our daily lives through devices such as smartphones [3] and autonomous vehicles [4]. However, the computational complexity and resource demands of deep neural networks (DNNs), typically trained and executed in high-precision floating-point arithmetic (e.g., 32-bit FP32), pose significant challenges for deployment in resource-constrained environments [5]. Neural network quantization has emerged as a pivotal technique to address these challenges by converting high-precision floating-point weights and activations into lower-precision formats, reducing memory usage, computational demands, and energy consumption with minimal accuracy loss [6]. This technique enables DNNs to operate effectively on resource-limited edge devices such as smartphones [7].
However, training quantized neural networks presents unique challenges despite their efficiency in inference. Unlike conventional deep neural networks, which rely on gradient-based optimization in a continuous domain, quantized models require solving a discrete optimization problem. The quantization process restricts the weights and activations to a predefined set of discrete values, resulting in a combinatorial search space and landscape [8]. To overcome these issues, studies on quantized neural network training focused primarily on two approaches: post-training quantization (PTQ) and quantization-aware training (QAT) [8]. PTQ transforms a pre-trained full-precision model into a low-bit representation by applying techniques such as range calibration or adaptive rounding, exemplified by AdaRound [9]. This method is computationally efficient but tends to suffer from significant accuracy loss under aggressive quantization, particularly at 4-bit precision or lower. In contrast, QAT integrates quantization effects during training, leveraging simulated quantization operations and the straight-through estimator (STE) to improve accuracy in low-precision scenarios [10]. Despite its advantages, QAT relies on heuristic approximations to circumvent the challenges posed by discrete optimization, limiting its ability to fully exploit the combinatorial structure of the problem.
These limitations highlight the need for a different approach to training quantized neural networks—one capable of efficiently navigating the discrete optimization landscape. The Ising machine [11, 12, 13, 14, 15, 16, 17] offers a promising avenue to address this challenge. By leveraging the energy-minimization properties of Ising models, researchers have applied Ising machines to train neural networks and perform statistical sampling [18]. For example, photonic Ising machines have demonstrated capabilities in low-rank combinatorial optimization and statistical learning tasks [19]; Equilibrium propagation using Ising machines has been utilized as a biologically plausible alternative to backpropagation for training neural networks effectively [20], and sparse Ising machines have been employed to train deep Boltzmann networks [21]. Using the principles of quantum coherence and Ising spin dynamics, these systems offer a unique approach to efficiently solving optimization problems central to neural network training, as well as in other fields such as wireless communication and signal processing [22, 23], molecular docking [24], computer vision [25, 26, 27, 28, 29, 30], etc.
In neural network training, the selection of the activation function critically influences the performance of the model and the adaptability to the task [31]. While the simplicity and computational efficiency of Rectified Linear Units (ReLUs) make them suitable for many applications, certain scenarios demand the distinctive non-linear characteristics of sigmoid or hyperbolic tangent (tanh) functions. Restricting the network to a specific activation function limits its expressive power and adaptability, potentially leading to suboptimal performance on certain tasks. Consequently, the development of training methodologies capable of supporting diverse activation functions becomes essential, as this versatility allows neural networks to effectively address broader problem domains and dataset characteristics, thereby significantly enhancing their practical applicability. A pioneering effort in this domain is the complete quantum neural network (CQNN) framework [32], which implements a comprehensive neural network architecture encompassing weights, biases, activation functions, and loss minimization within a quantum annealing system. This approach achieves the training in a single annealing step, offering notable advantages, including guaranteed convergence to global optima and substantially reduced training durations, in stark contrast to the iterative optimization process characteristic of conventional gradient descent methods.
However, a critical limitation of CQNN emerges in its handling of activation functions. To conform to the quadratic constraints of the Ising model Hamiltonian, CQNN approximates arbitrary activation functions (e.g., ReLU, sigmoid) with polynomials. This choice raises concerns rooted in approximation theory. As established by Leshno et al. [33], multilayer feedforward networks with non-polynomial activation functions are dense in the space of continuous functions between Euclidean spaces under the compact convergence topology, a property ensuring universal approximation capability. Polynomial activation functions, conversely, limit this expressiveness, as they form a finite-degree polynomial space incapable of approximating arbitrary continuous functions [33, 34]. In addition to the limitations in expressive power, another significant drawback that arises when constructing QUBO models is, polynomial activation functions require a large number of Ising spins for degree reduction. This issue further exacerbates the challenges of applying CQNN in practical scenarios. Therefore, adapting arbitrary, non-polynomial activation functions to quantum computing frameworks remains an open challenge, requiring innovative encoding strategies or hybrid quantum-classical solutions to preserve both computational efficiency and model versatility.
When addressing constrained quadratic optimization problems, the penalty method is frequently used to incorporate constraints into the objective function [35]. However, this approach has significant limitations. First, quadratic penalty functions introduce pairwise interactions among all variables, transforming sparse problems into fully connected ones. This increased connectivity requires complex mapping techniques, particularly for quantum annealers with limited connectivity [36], substantially raising resource demands on near-term quantum hardware [37]. Second, quadratic penalties generate large energy scales that can exceed hardware dynamic ranges, necessitating Hamiltonian normalization [37]. This normalization reduces the problem’s effective energy resolution, impairing optimization performance. Additionally, the penalty method alters the optimization landscape, requiring careful tuning of penalty coefficients, a critical hyperparameter for effective constraint handling. As the number of constraints grows, the search space for optimal penalty coefficients expands exponentially, posing a significant challenge to achieving robust optimization outcomes.
In this study, we present a framework for training quantized neural networks using CIM.
Our main contributions are summarized as follows.
  • Unified quantum-compatible and convex formulation for neural networks training. We develop a general spline-based discretization scheme combined with Forward Interval Propagation (FIP), which transforms arbitrary activation and loss functions into a QCBO model. By leveraging copositive programming, we show that this QCBO formulation admits an exact convex lifting, allowing (quantized and full precision) neural networks of arbitrary depth, width, and piecewise polynomial activations to be represented as a single convex optimization problem over the completely positive cone. The resulting CP formulation preserves all nonlinear and combinatorial interactions exactly, while maintaining a convex feasible region in the lifted space. Moreover, we derive a sample complexity bound from an optimization perspective, effectively constraining the number of spins required in the quadratic binary model. This bound lays a critical foundation for scaling to large-scale scenarios by ensuring computational feasibility and efficiency.
  • A scalable quantum classical optimization framework with provable robustness. We introduce the noise-robust version of Quantum Conditional Gradient Descent (QCGD) algorithm tailored to the copositive formulation and prove its convergence under realistic quantum oracles with randomness and limited coefficient precision. To further address the scalability challenge of the monolithic lifted model, we adopt a Decomposed Lower-Bound Optimization (DLBO) strategy that replaces the global cone with sample-wise copositive blocks [38, 39]. This decomposition enables efficient bit-level optimization per sample and plays a role analogous to stochastic gradient descent in classical deep learning. It substantially reduces the dimension of each quantum subproblem while preserving the tightness of the relaxation. We further show how both the QCGD algorithm and newly introduced Quantum Progressive Hedging (QPH) algorithm can be adopted to solve the decomposed problem.
  • Training and empirical validation. Our experimental evaluation highlights the effectiveness of this approach across multiple dimensions. Compared to standard quantization-aware training (QAT) methods, our method achieves higher accuracy with training times reduced by several orders of magnitude. When evaluated against post-training quantization (PTQ) algorithms, it delivers superior accuracy with fewer parameter bits, achieving $94.95\%$ accuracy with only $1.1$-bit precision. Furthermore, compared to full-precision models with continuous weights, our quantized networks significantly reduce memory footprint and inference latency while maintaining competitive accuracy. To further validate the scalability and effectiveness of the proposed DLBO-based training framework, we extend our evaluation beyond binary classification and conduct experiments on multiclass classification tasks using the Fashion-MNIST, Wine, and Digits datasets. Across all datasets and task settings, our method consistently delivers high accuracy.
These results demonstrate the potential of quantum optimization to enable efficient and high-performing deep learning models for edge deployment and pave the way for the next generation of efficient deep learning systems.
Results
Principle
The approach first conceptualizes a quantized neural network (QNN) as a parameterized operator $\mathcal{Q}:\mathbb{R}^{d}\to\mathbb{R}$, which transforms an input space $\mathbb{R}^{d}$ into the output space $\mathbb{R}$ through a parameter set $\mathcal{W}$ drawn from a discrete parameter space. This operator emerges from a hierarchical cascade of quantized transformations, reflecting the network’s layered and constrained architecture.
The global operator $\mathcal{Q}(x,\mathcal{W})$ could be expressed as a nested sequence of $L$ quantized layers:
$\mathcal{Q}(x,\mathcal{W})=\mathcal{Q}_{L}(\mathcal{Q}_{L-1}(\cdots\mathcal{Q}_{1}(x,\mathcal{W}_{1}),\mathcal{W}_{2})\cdots,\mathcal{W}_{L}),$ (1)
where $L$ denotes the depth of the network, and $\mathcal{W}$ encompasses the quantized parameters across all layers. For each layer $l$ ($1\leq i\leq L$), the parameter set $\mathcal{W}_{l}=(\mathbf{W}_{l},\mathbf{b}_{l})$ consists of a quantized weight matrix $\mathbf{W}_{l}$ and a quantized bias vector $\mathbf{b}_{l}$, with dimensional compatibility where the input dimension at layer $1$ is $d$ and the output dimension at layer $L$ is 1.
Each layer’s transformation $\mathcal{Q}_{l}$ operates on the quantized output of the previous layer via the transform as:
$\mathcal{Q}_{l}(x,\mathcal{W}_{l})=\sigma(\mathbf{W}_{l}\cdot x+\mathbf{b}_{l}),$ (2)
where $\sigma:\mathbb{R}\to\mathbb{R}$ denotes the non-linear activation function applied elementwise, and the operation $\cdot$ reflects a quantized linear combination.
The activation value of the $j^{\text{th}}$ neuron in the $l^{\text{th}}$ layer is related to the activation values of the $(l-1)^{\text{th}}$ layer through the following formula:
$a_{l,j}=\sigma\Big(\sum_{k}{w_{ljk}a_{l-1,k}+b_{lj}}\Big).$ (3)
Here, the sum of $k$ is taken over all the neurons in the (l1)(l-1)-th layer.
Figure 1: The flowchart of the quantized neural network training process using CIM and hybrid techniques.
Figure 1 illustrates the workflow of the quantized neural network training framework, using the hybrid approach and CIM. The process begins with data preparation and progresses through the formulation of the loss landscape and solution space of a hypercube structure. Key steps include the constraint linearization, the transformation of the problem into a QCBO model, and the iterative solution process using an Ising machine. Variables are updated based on the QUBO matrix at each iteration, ensuring convergence toward the optimal solution. Therefore, this work can achieve marked performance in quantum computing in real-world applications, laying a solid foundation for using quantum computing in QNN training.
Spline-Induced Neural Network and Optimization Landscape
The piecewise linear functions are adopted as the activation functions in the neural network. These functions are particularly advantageous due to their simplicity and computational efficiency, while still maintaining the ability to approximate complex, non-linear relationships, as shown in the following fact.
Fact 1 ([40, 41])
For any function $f\in\mathcal{F}_{d,n}=\mathcal{W}^{n,\infty}([0,1]^{d})$ (where $d$ is the input dimension, $n$ is the order of weak derivatives, and $\max_{\mathbf{n}:|\mathbf{n}|\leq n}\text{ess sup}_{\mathbf{x}\in[0,1]^{d}}|D^{\mathbf{n}}f(\mathbf{x})|\leq 1$), and any error bound $\epsilon\in(0,1)$, there exists a quantized neural network $f_{q}$ with any piecewise linear activation function having a finite number of breakpoints and weights restricted to $\lambda\geq 2$ discrete values, such that:
$|f(\mathbf{x})-f_{q}(\mathbf{x})|\leq\epsilon,\quad\forall\mathbf{x}\in[0,1]^{d}.$ (4)
The number of weights required by $f_{q}$ has an upper bound of $\mathcal{O}(\lambda\log^{\frac{1}{\lambda-1}+1}(1/\epsilon)(1/\epsilon)^{d/n})$, while an unquantized piecewise linear neural network $f_{u}$ with the same activation function achieving the same error has an upper bound of $\mathcal{O}(\log(1/\epsilon)(1/\epsilon)^{d/n})$. Thus, the quantized weight count exceeds the unquantized one by a factor of $\mathcal{O}(\log^{5}(1/\epsilon))$, which is significantly smaller than the lower bound for the unquantized network, $\Omega(\log^{-3}(1/\epsilon)(1/\epsilon)^{d/n})$.
This fact confirms that a quantized piecewise linear neural network, regardless of the specific activation, approximates target functions with efficiency comparable to an unquantized one, incurring only a low complexity overhead. Ding et al. [40] further reveal a theoretical optimal number of discrete weight values, which typically lies between 1 and 4 bits across a wide range of $d$ and $\epsilon$, indicating that effective quantization of piecewise linear neural networks requires only a small number of discrete weights. The efficiency supports the use of quantized piecewise linear neural networks in quantum computing contexts, where minimal discrete states could leverage quantum parallelism for training.
While Fact 1 establishes the universal approximation capability of quantized piecewise linear neural networks, it remains an interesting question to quantify the approximation error introduced when replacing a smooth activation function with a piecewise linear one. For completeness, we provide the error bound induced by piecewise linear approximations in a feedforward neural network in Section 2 of the supplementary material.
The objective of training can be formulated as a discrete optimization problem over the parameter space, aiming to minimize a generalized loss function evaluated across the entire input distribution $x$:
$\min_{\mathcal{W}\in\mathcal{W}_{\text{quant}}}\sum_{x}\mathcal{L}(a^{L}_{x},y_{x}),$ (5)
subject to the quantized layer dynamics:
$a^{l}_{x}$ $=\sigma(z^{l}_{x}),\quad\forall x,\;1\leq l\leq L,$ (6)
$z^{l}_{x}$ $=\mathbf{W}_{l}\cdot a^{l-1}_{x}+\mathbf{b}_{l},\quad\forall x,\;1\leq l\leq L,$ (7)
where $a^{0}_{x}=x$ serves as the input, $a^{L}_{x}$ is the quantized network output, $y_{x}$ represents the target output, and $\mathcal{W}_{\text{quant}}$ denotes the set of all permissible quantized parameter configurations. This formulation captures the trade-offs between accuracy and computational efficiency inherent in quantization.
Forward Interval Propagation.
Activation functions play a crucial role in neural networks by introducing nonlinearities that enable complex mappings between inputs and outputs. We leverage spline approximation with piecewise constant functions to approximate activation and loss functions, such as mean squared error (MSE), cross-entropy loss, and Huber loss, which quantify the difference between predicted and actual values in machine learning tasks.
Forward Interval Propagation(FIP) is designed to address the challenges of non-linearity in neural networks, particularly the complex multi-layer composite relationships inherent in activation functions. Also, FIP is to simplify these complex interactions by discretizing the activation function into multiple linear subintervals. This allows for efficient computation while preserving the expressive power of neural networks. As shown in Figure 2, we divide the entire value range into multiple subintervals. The activation function is defined as a piecewise function, where the output of each neuron is determined by the specific subinterval it belongs to. The summation of these outputs, along with the bias term, determines the subinterval for the output of the next layer. This process is repeated during forward propagation, establishing a relationship between the input and output by propagating through the defined subintervals.
One of the key benefits of FIP is its ability to capture and manage the multi-layer composite relationships in neural networks. As the forward propagation proceeds, the activation functions at each layer are constrained to a specific subinterval, and the overall relationship between input and output is effectively captured by the combination of these subintervals.
Figure 2: (a) Feedforward neural network with piecewise linear activation function. (b) Schematic representation of forward interval propagation (FIP) in a neural network, illustrating the discretization of activation functions into linear subintervals. The diagram shows multiple input neurons (left) connected through weights $(w_{0},w_{1},\ldots,w_{m})$ and a bias term $b$ to a summation node $\Sigma$, followed by an activation function. The green-highlighted intervals and values of $\beta_{i}$s indicate the specific subintervals that the activation function’s output values belong to, determining the output subinterval for the next layer, effectively capturing multi-layer composite relationships during forward propagation.
We divide the interval $\underline{M},\overline{M}]$ as $\underline{M}=M_{0}\leq M_{1}\leq M_{2}\leq\ldots\leq M_{n}=\overline{M},$ and the neural network training problem could be formulated as the following QCBO expression:
$\min$ $\sum_{x}\sum_{i=1}^{n+1}\beta_{a^{L}_{x}}^{(i-1)}\cdot\mathcal{L}(M_{i-1},y_{x})$ (8)
s.t. $a^{l}_{x,j}=\sum_{i=1}^{n+1}\beta_{z^{l}_{x,j}}^{(i-1)}\cdot\frac{\sigma(M_{i-1})+\sigma(M_{i})}{2},\quad\forall x,j,l,$ (9)
$\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}M_{i-1}\leq a^{L}_{x}\leq\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}M_{i},\quad\forall x,$ (10)
$\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}M_{i-1}\leq\sum_{k}(w^{l}_{jk}a^{l-1}_{x,k}+b^{l}_{j})\leq\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}M_{i},\quad\forall x,j,l,$ (11)
$\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}=1,\quad\forall x,$ (12)
$\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}=1,\quad\forall x,j,l,$ (13)
where $x$, $y_{x}$ represents the label, $a^{L}_{x}$ denotes the output of the last layer; $z^{l}_{x,j}$ and $a^{l}_{x,j}$ denotes the input and output of the $j$-th neuron of the $l$-th layer; $\mathcal{L}(M_{i},y_{x})$ indicates the loss when the last layer’s output is in the $i$-th subinterval and the target is $y_{x}$. For any variable $\tau$, $\beta^{(i)}_{\tau}=1$ if $\tau$ lies within the $i$-th sub-interval and 0 otherwise.
Note that $w^{l}_{jk}a^{l-1}_{x,k}$ contains quadratic terms, which will introduce higher-order terms when transformed into the QUBO form. Therefore, it is necessary to perform a reduction operation to the lower order. For example, when using the penalty method to obtain the QUBO model, we can replace the term $x_{1}x_{2}$ by a binary variable $y$ and add the Rosenberg polynomial [42] $h(x_{1},x_{2},y)=3y+x_{1}x_{2}-2y(x_{1}+x_{2}),$ to the objective function. These auxiliary variables help to linearize the expression, making it easier to handle.
In addition, the non-binary variables $a^{l}_{x,j}$, $w^{l}_{jk}$, $b^{l}_{j}$ are represented using binary encoding as following: $a^{l}_{x,j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},w^{l}_{jk}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{w^{l}_{jk}},b^{l}_{j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{b^{l}_{j}}$. In Section 3 of supplementary, we provide further details of spline approximation.
From Multi-level Quantization to Binary [43].
For layer $\ell\in[L]$, let the trainable weight tensor be vectorized as $\mathbf{w}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ and each coordinate $p\in[d_{\ell}]$ be constrained to a finite codebook
$w^{(\ell)}_{p}\ \in\ \mathcal{S}^{(\ell)}_{p}\ \subset\ \mathbb{R},\qquad\mathcal{S}^{(\ell)}_{p}=\{\alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,1}+\beta^{(\ell)}_{p},\ \dots,\ \alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,K_{p}}+\beta^{(\ell)}_{p}\}$, (14)
where $\alpha^{(\ell)}_{p},\beta^{(\ell)}_{p}\in\mathbb{R}$ allow per-layer or per-channel scaling/shift and $\{c^{(\ell)}_{p,k}\}_{k=1}^{K_{p}}$ are fixed codes (not necessarily equally spaced or symmetric). Biases are treated identically.
One-hot encoding of weights. Introduce for each $(\ell,p)$ the binary indicator block
$\mathbf{y}^{(\ell)}_{p}=(y^{(\ell)}_{p,1},\dots,y^{(\ell)}_{p,K_{p}})^{\top}\in\{0,1\}^{K_{p}},\qquad\sum_{k=1}^{K_{p}}y^{(\ell)}_{p,k}=1,$ (15)
and define the affine map
$w^{(\ell)}_{p}\ =\ \sum_{k=1}^{K_{p}}\Big(\alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,k}+\beta^{(\ell)}_{p}\Big)\,y^{(\ell)}_{p,k}$. (16)
Table 1 summarizes the required number of Ising spins for various variables in our quadratic programming model, showing that the overall number of Ising spins required is $O(nWLN)$, for handling the datasets with $N$ samples and the neural network with width $W$ and depth $L$.
Variable Required Number of Ising spins
$\beta^{(i)}_{x}$ $O(nWLN)$
$w^{l}_{jk}$ $O(LW^{2})$
$b^{l}_{j}$ $O(WL)$
$a^{l}_{x,j}$ $O(WLN)$
Linearization $O(nWLN)$
Table 1: The number of required Ising spins for each variable.
Recall that our objective is to minimize:
$\mathcal{L}(\theta)=\mathbb{E}_{(x,y_{x})\sim\mathcal{P}}[C_{x}(f_{\theta}(x),y_{x})],$ (17)
where $\theta=\{w^{l},b^{l}\}_{l=1}^{L}$, $f_{\theta}(x)=a^{L}_{x}$. Then, we solve the empirical risk minimization (ERM) over $N$ i.i.d. samples $\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{P}$:
$\hat{\theta}_{N}=\arg\min_{\theta\in\Theta}\hat{\mathcal{L}}_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C_{x_{i}}(f_{\theta}(x_{i}),y_{i}).$ (18)
Note that the number of samples $N$ can be bounded while ensuring the efficiency of our learning process; this is where the concept of VC dimension becomes crucial.
Lemma 1 ([44, 45])
The VC dimension of a neural network with piecewise polynomial activation functions is bounded by $O(W^{3}L^{2})$. Specifically, when the activation functions are piecewise constant, the VC dimension is no more than $O(W^{2}L(\log W+\log L))$.
Within the Probably Approximately Correct (PAC) learning framework [46], we focus on controlling the generalization error, a measure of how accurately an algorithm is able to predict outcomes for previously unseen data. This framework ensures that, with high probability $1-\alpha$, the generalization error is confined within a precision $\varepsilon$. Consequently, the required number of samples can be bounded as:
$N=O\Big(\frac{\text{VCdim}}{\varepsilon}\log\frac{1}{\alpha}\Big)=O\Big(\frac{W^{2}L(\log W+\log L)}{\varepsilon}\log\frac{1}{\alpha}\Big).$ (19)
Moreover, from an optimization perspective, we can derive a sample complexity bound tailored to the optimization error, defined as the discrepancy between the expected loss of the learned parameters and the minimum loss, i.e., $\mathcal{L}(\hat{\theta}_{N})-\mathcal{L}(\theta^{*})$. This focus on optimization error is particularly pertinent given that our study is about addressing optimization problems,
Theorem 1
For a quantized neural network with $B$-bit parameters, let $\varepsilon,\alpha\in(0,1)$, and assume the loss function $\mathcal{L}(\cdot)$ is bounded in $0,\mathcal{L}_{\max}]$. Then, with probability at least $1-\alpha$,
$\mathcal{L}(\hat{\theta}_{N})\leq\mathcal{L}(\theta^{*})+\varepsilon,$ (20)
provided $N=O\left(\frac{\mathcal{L}_{\max}^{2}}{\varepsilon^{2}}\log\frac{|\Theta|}{\alpha}\right)$. Substituting $|\Theta|$:
$N=O\left(\frac{\mathcal{L}_{\max}^{2}}{\varepsilon^{2}}\sum_{l=1}^{L}d_{l}(d_{l-1}+1)B+\log\frac{1}{\alpha}\right),$ (21)
where $d_{l}$ and $d_{l-1}$ are the layer dimensions.
Proof: Define the function class $\mathcal{G}=\{g_{\theta}(x,y)=\mathcal{L}(f_{\theta}(x),y):\theta\in\Theta\}$. Given the finite $\Theta$, apply Hoeffding’s inequality for $g_{\theta}\in[0,\mathcal{L}_{\max}]$:
$P\left(\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right).$ (22)
Using a union bound over all $|\Theta|$ functions:
$P\left(\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right).$ (23)
Set the probability to $\alpha$, i.e., $2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right)\leq\alpha$, we have $N\geq\frac{\mathcal{L}_{\max}^{2}}{2\varepsilon^{2}}\log\frac{2|\Theta|}{\alpha}$. Since $\hat{\theta}_{N}$ minimizes $\hat{\mathcal{L}}_{N}$:
$\mathcal{L}(\hat{\theta}_{N})\leq\hat{\mathcal{L}}_{N}(\hat{\theta}_{N})+\varepsilon/2\leq\hat{\mathcal{L}}_{N}(\theta^{*})+\varepsilon/2\leq\mathcal{L}(\theta^{*})+\varepsilon.$ (24)
Since the total number of configurations is $|\Theta|\leq\prod_{l=1}^{L}(2^{B})^{d_{l}(d_{l-1}+1)}$. The proof is complete by substituting $|\Theta|$.
$\Box$
Convex Formulation for Neural Network Training
For each sample $s$ and layer $\ell$, we denote pre-activations and post-activations by $\mathbf{z}^{(s,\ell)}\in\mathbb{R}^{m_{\ell}}$ and $\mathbf{a}^{(s,\ell)}\in\mathbb{R}^{m_{\ell}}$ with $\mathbf{a}^{(s,0)}=\mathbf{x}^{(s)}$. All decision variables are stacked in a single vector
$\mathbf{u}=\big[\,\bm{\delta}^{(1{:}L)}_{W},\ \bm{\delta}^{(1{:}L)}_{b},\ \{\mathbf{v}^{(s,\ell)}\}_{s,\ell},\ \{\mathbf{u}^{(s,\ell)}_{\!\mathrm{bil}}\}_{s,\ell},\ \{\mathbf{z}^{(s,\ell)},\mathbf{a}^{(s,\ell)}\}_{s,\ell},\ \{\bm{\beta}^{(s,\ell)}_{z}\}_{s,\ell},\ \{\bm{\beta}^{(s)}_{a}\}_{s},\ \mathbf{s}\,\big]^{\top},$ (25)
where $\bm{\delta}$ are binary code bits, $\mathbf{v}$ are auxiliary bitwise products, $\mathbf{u}_{\!\mathrm{bil}}$ are summed bilinear terms, $\bm{\beta}$ are one-hot selectors, and $\mathbf{s}\geq 0$ collects inequality slacks.
We adopt the standard CP lifting
$\mathbf{X}:=\begin{bmatrix}\bm{\Lambda}&\mathbf{u}\\ \mathbf{u}^{\top}&1\end{bmatrix}\in\mathcal{C}^{\ast},$ (26)
with $\mathcal{C}^{\ast}$ the cone of completely positive matrices. Every linear row imposed on $\mathbf{u}$ is paired with its self-quadratic mate in $\bm{\Lambda}$. The quantized weights and biases are generated from binary codes by affine codebooks:
$\operatorname{vec}\!\big(\mathbf{W}^{(\ell)}\big)$ $=\mathbf{C}^{(\ell)}_{W}\,\bm{\delta}^{(\ell)}_{W}+\mathbf{d}^{(\ell)}_{W},\qquad\mathbf{b}^{(\ell)}=\mathbf{C}^{(\ell)}_{b}\,\bm{\delta}^{(\ell)}_{b}+\mathbf{d}^{(\ell)}_{b}$. (27)
Stacking $\ell=1$ to $L$ yields block diagonal $\mathbf{C}_{W}=\mathrm{diag}(\mathbf{C}^{(1)}_{W},\dots,\mathbf{C}^{(L)}_{W})$ and $\mathbf{C}_{b}$.
For the quadratice terms appear in (11), we introduce the following linearization procedure. Let the propagated activation bounds satisfy $\underline{\mathbf{a}}^{(s,\ell-1)}\leq\mathbf{a}^{(s,\ell-1)}\leq\overline{\mathbf{a}}^{(s,\ell-1)}$ componentwise. For each edge $(j,k)$ in layer $\ell$ and each bit $r$ of $W^{(\ell)}_{jk}$, we introduce an auxiliary continuous variable
$v^{(s,\ell)}_{jk,r}\;=\;\delta^{(\ell)}_{jk,r}\,a^{(s,\ell-1)}_{k}$, (28)
where $\delta^{(\ell)}_{jk,r}\in\{0,1\}$ is the binary code bit for that weight component, and $a^{(s,\ell-1)}_{k}$ is a bounded continuous activation. We enforce this product using the standard four McCormick linear inequalities (valid for $\delta\in\{0,1\}$, $a\in[\underline{a},\overline{a}]$):
$v^{(s,\ell)}_{jk,r}\;\leq\;\overline{a}^{(s,\ell-1)}_{k}\,\delta^{(\ell)}_{jk,r}$, $\text{(upper envelope at $a=\overline{a}$)},$
$v^{(s,\ell)}_{jk,r}\;\geq\;\underline{a}^{(s,\ell-1)}_{k}\,\delta^{(\ell)}_{jk,r}$, $\text{(lower envelope at $a=\underline{a}$)},$
[…]
Based on the defintions and linearization above, we present the details of constructing the linear constraints in Section 4 of the supplementary. Then, we collect all equalities $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q}$ such as pre-activation and post-activation constraints, and all inequalities $\mathbf{N}_{r}\mathbf{u}\leq\mathbf{n}_{r}$ including McCormick and interval constraints, etc.), then turn each inequality into an equality with slack $\mathbf{s}_{r}\geq 0$.
According to [47], we add the following self-quadratic constraint,
$\langle\mathbf{M}_{q}^{\top}\mathbf{M}_{q},\bm{\Lambda}\rangle=\|\mathbf{m}_{q}\|_{2}^{2},\qquad\langle\mathbf{N}_{r}^{\top}\mathbf{N}_{r},\bm{\Lambda}\rangle+2\,\mathbf{N}_{r}\,\bm{\Lambda}[\mathbf{u},\mathbf{s}_{r}]+\bm{\Lambda}[\mathbf{s}_{r},\mathbf{s}_{r}]=\|\mathbf{n}_{r}\|_{2}^{2}.$ (30)
The empirical risk is linear in $\mathbf{u}$, i.e., the objective can be expressed as
$\min\ \langle\mathbf{Q}_{0},\mathbf{X}\rangle,\qquad\mathbf{Q}_{0}=\begin{bmatrix}\mathbf{0}&\tfrac{1}{2}\mathbf{c}^{\top}\\ \tfrac{1}{2}\mathbf{c}&0\end{bmatrix}$, (31)
and the full completely positive program (CPP) is as following,
$\min_{\mathbf{X}}\quad$ $\langle\mathbf{Q}_{0},\mathbf{X}\rangle$ (32)
s.t. $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q},\ \ \langle\mathbf{M}_{q}^{\top}\mathbf{M}_{q},\bm{\Lambda}\rangle=\|\mathbf{m}_{q}\|_{2}^{2},\quad\forall q,$ (33)
$\mathbf{N}_{r}\mathbf{u}+\mathbf{s}_{r}=\mathbf{n}_{r},\ \ \langle\mathbf{N}_{r}^{\top}\mathbf{N}_{r},\bm{\Lambda}\rangle+2\,\mathbf{N}_{r}\bm{\Lambda}[\mathbf{u},\mathbf{s}_{r}]+\bm{\Lambda}[\mathbf{s}_{r},\mathbf{s}_{r}]=\|\mathbf{n}_{r}\|_{2}^{2},\ \forall r,$ (34)
$\begin{bmatrix}\bm{\Lambda}&\mathbf{u}\\ \mathbf{u}^{\top}&1\end{bmatrix}\in\mathcal{C}^{\ast}.$ (35)
While the above formulation assumes trainable parameters are encoded as binary vectors, which leads to mixed binary-continuous products that we linearize via McCormick inequalities before lifting to a completely positive program, the same convex lifting paradigm can be extended to continuous weights and biases. Furthermore, it also generalizes to networks with piecewise-polynomial activation functions.
As shown in (8)-(13), the problem of training a full precision neural network can be written as a quadratically constrained quadratic program (QCQP). It is known that such QCQP can be represented exactly as an optimization problem with a linear objective and constraints over a generalized copositive cone of the form [48]
$\mathcal{C}^{\ast}(\mathcal{K})\ :=\ \Big\{\,\mathbf{X}\in\mathbb{S}^{p+1}_{+}\;\Big|\;\mathbf{X}=\sum_{r}\begin{bmatrix}\mathbf{y}^{(r)}\\ 1\end{bmatrix}\begin{bmatrix}\mathbf{y}^{(r)}\\ 1\end{bmatrix}^{\!\top},\quad\mathbf{y}^{(r)}\in\mathcal{K}\ \text{for all }r\Big\},$ (36)
where $\mathcal{K}$ is any closed convex cone encoding the domain restrictions, and $\mathbb{S}^{p+1}_{+}$ denotes symmetric $(p{+}1)\times(p{+}1)$ matrices. [^1] Classical copositive programs correspond to the special case $\mathcal{K}=\mathbb{R}^{p}_{+}$, allowing a general convex $\mathcal{K}$ yields a generalized copositive cone.
The same mechanism extends beyond piecewise linear activations. Suppose each scalar activation is given by a piecewise polynomial of fixed degree $d$, e.g. on each interval $M_{i-1},M_{i}]$,
$a=p_{i}(z),\qquad p_{i}(z)=\sum_{r=0}^{d}\gamma_{i,r}\,z^{r},$ (37)
together with one-hot interval selection. We can lift these polynomial relations to at most quadratic form by introducing auxiliary monomial variables. Any monomial with degree larger than two can reduced to quadratic form by recursively introducing auxiliary continuous variables, i.e., for a cubic term $z_{1}z_{2}z_{3}$, we introduce $y_{12}=z_{1}z_{2}$ and rewrite $z_{1}z_{2}z_{3}$ as $y_{12}z_{3}$. After this recursive degree reduction, the entire network with continuous weights and piecewise-polynomial activations remains a QCQP, and thus still admits an exact convex reformulation via the generalized copositive lifting described above.
Geometrically, a single globally optimal CP solution $V^{*}$ corresponds to a whole family of parameterizations in the original weight space, permutations as well as redundancies in the binary encoding, all realize the same $V^{*}$.
Noise-Robust Hybrid Dynamics
In this section, we confront the formidable optimization challenges stemming from the proliferation of constraints in large-scale networks in the aforementioned model. The abundance of constraints not only complicates the optimization process but also necessitates the determination of numerous penalty coefficients in the traditional penalty method. The task of determining these penalty coefficients becomes increasingly complex as the number of constraints increases, leading to potential inefficiencies and suboptimal solutions.
To surmount this hurdle, we use the Quantum Conditional Gradient Descent (QCGD) algorithm developed in [49], a cutting-edge hybrid classical-quantum framework tailored to address quadratic, linearly-constrained, binary optimization problems on quantum machines. Our workflow starts with data being fed into the neural network. The optimization problem is then formulated as a QCBO. This problem is subsequently linearized and transformed into a copositive programming problem. In the hybrid quantum-classical iterative process, during each iteration, the quantum conditional gradient method calculates the update direction by minimizing a linear approximation of the penalized proxy of the objective function. QCGD formulates the linear minimization problem as a QUBO sub-problem, which is suitable for Ising machines. CIM solves the QUBO problem, and the solution obtained is fed back into the classical computing step to calculate the direction of iteration. This iterative loop continues, with the QUBO matrix being updated at each step, until convergence towards an optimal solution is achieved.
In the following theorem, we prove that by sampling a certain number of solutions from the Ising machine at each iteration, the algorithm can still converge despite the presence of errors.
Theorem 2
Consider the Quadratic Constrained Binary Optimization (QCBO) problem, which could be solved using the QCGD algorithm with a random quantum oracle as defined in Definition 1. At step $t$, the error in the objective in a single run of the quantum oracle has mean $\mu_{t}$ and variance $\sigma_{t}$, both of which have time-independent upper bounds. Let $T$ denote the total number of iterations of the QCGD algorithm, then:
  1. 1.
    QCGD algorithm converges to the optimal solution almost surely, i.e., let $V_{t}$ denote the solution matrix at iteration $t$ generated by the QCGD algorithm, and let $V^{*}$ denote the optimal solution of the QCBO problem in the completely positive cone, then
    $\mathbb{P}(\lim_{t\rightarrow\infty}\mathbf{V}_{t}=\mathbf{V}^{*})=1.$ (38)
  2. 2.
    In expectation, the objective gap and residual gap converge as:
    $\mathbb{E}[\text{Objective-gap}_{T}]=O\Big(\frac{1}{\sqrt{T}}\Big),\;\mathbb{E}[\text{Infeasibility}_{T}]=O\Big(\frac{1}{\sqrt{T}}\Big)$ (39)
  3. 3.
    The total time to solution for any fixed success probability satisfies:
    $\text{TTS}_{\text{QCBO}}=O\left(\tau\cdot T\cdot\log n\right),$ (40)
where
  • $m_{t}=c\log T+1$ is the number of independent quantum samples per iteration $t$, with $c=\frac{2}{-\log(1-p_{0})}$, $p_{0}>0$ the probability of the quantum oracle returning the optimal QUBO solution,
  • $\tau$ is the time per quantum oracle call.
Proof: The subproblem in the $t$-th iteration is solved via $m_{t}=c\log n$ repetitions, selecting the best solution among samples $\{\mathbf{D}^{(i)}_{t}\}_{1\leq i\leq m_{t}}$, i.e., $\mathbf{D}_{t}=\mathbf{D}^{(i_{t})}_{t}$, where
$i_{t}=\operatorname*{argmin}_{1\leq i\leq m_{t}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{(i)}_{t})$, (41)
then $\mathbf{V}_{t+1}=(1-\eta_{t})\mathbf{V}_{t}+\eta_{t}\mathbf{D}_{t}$. For any $r>0$, the $r$-th moment of quantum oracle error
$\mathbb{E}\Big[\Big(\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big)^{r}\Big]$ (42)
$=$ $\mathbb{P}\Big(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big)\cdot\mathbb{E}\Big[\Big(\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big)^{r}\Big|\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big]$ (43)
$\leq$ $\mathbb{P}\Big(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big)\cdot\mathbb{E}\Big[\Big(\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{(i)}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big)^{r}\Big|\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big]$ (44)
$=$ $\mathbb{P}\Big(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big)\cdot\mathbb{E}\Big[\Big(\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{(i)}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big)^{r}\Big|\mathbf{D}^{(i)}_{t}\neq\mathbf{D}^{*}_{t}\Big]$ (45)
$=$ $\frac{\mathbb{P}\Big(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t}\Big)}{\mathbb{P}\Big(\mathbf{D}^{(i)}_{t}\neq\mathbf{D}^{*}_{t}\Big)}\cdot\mathbb{E}\Big[\Big(\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{(i)}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big)^{r}\Big],$ (46)
where the second equality is based on the independence of samples. For $m_{t}=c\log T$,
$\frac{\mathbb{P}(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t})}{\mathbb{P}(\mathbf{D}^{(i)}_{t}\neq\mathbf{D}^{*}_{t})}=(1-p_{0})^{m_{t}-1}=\frac{1}{T^{2}},$ (47)
we have
$\mathbb{E}\Big[\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big]$ $\leq\frac{\mu_{t}}{T^{2}},$ (48)
$\mathrm{Var}\Big[\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}_{t})-\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}^{*}_{t})\Big]$ $\leq\frac{\sigma_{t}+\mu^{2}_{t}}{T^{2}}.$ (49)
We next claim that $\mu_{t}$ and $\sigma_{t}$ is bounded, since
$\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D}_{t})\leq\sum_{i,j}{|Q_{ij}^{(t)}|}=O(n^{2}),$ (50)
which is independent of iteration index $t$, assuming each element in $\mathbf{Q}^{(t)}_{\text{QUBO}}$ is constant. According to Proposition 1, we can get the order of convergence in expectation, and the almost sure convergence follows from the concentration probability bound that converges to $1$ when $T$ is infinity.
The bound on TTS can also be derived from the concentration result in Proposition 1. A more straightforward approach is based on the union bound, the probability that
$\mathbb{P}(\exists\;t,\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t})\leq\sum_{t=1}^{T}{\mathbb{P}(\mathbf{D}_{t}\neq\mathbf{D}^{*}_{t})}=\frac{1}{T}$ (51)
Therefore, With high probability $1-\frac{1}{\mathrm{poly}(n)}\geq p_{R}$, $\mathbf{D}_{t}=\mathbf{D}^{*}_{t}$ for all $t$, yielding the convergence to optimal solution. As per-iteration time of QCGD is $m_{t}\tau$, we can obtain the second conclusion for $\text{TTS}_{\text{QCBO}}$. $\Box$
Quantum computers, such as those implementing adiabatic quantum computing (AQC) for solving QUBO problems, face challenges due to limited parameter precision [50]. Theoretically, QUBO parameters are real-valued; however, practical hardware—whether quantum annealers such as those from D-Wave or other accelerators—relies on finite-precision representations, typically using fixed-bit integers. This limitation introduces perturbations into the quantum system, distorting the energy landscape and potentially shifting the global optimum.
Considering directly truncating the Ising coefficients to $d$ digits, the following lemma shows that the convergence can be preserved under certain conditions.
Lemma 2
Let the coefficients of the QUBO problem be truncated to $d=O(\log n)$ digits. If QCGD is used with these truncated coefficients, it will retain convergence guarantees, and the number of iterations remains in the same order.
This result is appealing because it demonstrates that even with limited precision, the hybrid algorithm can still achieve convergence without a substantial increase in computational effort. This behavior can be interpreted as a form of implicit error correction. The introduced errors are effectively controlled and do not accumulate in a way that disrupts the convergence properties of the algorithm.
Figure 3: The monolithic copositive formulation constructs a single large cone whose dimension scales with the number of samples NN. DLBO replaces this with sample-wise complete positive cones that share global parameters $(\mathbf{x},\mathbf{X})$. Each epoch solves per-sample QUBO subproblems on quantum Ising hardware and aggregates their solutions to update the shared parameters, enabling scalable quantum–classical training.
Training with Single-Sample Bit-Scale Optimization
Training large-scale copositive programming formulations involves optimizing over the full parameter space across all $N$ samples, which can become computationally prohibitive in terms of the required number of Ising spins. To overcome this scalability challenge, we introduce a novel training paradigm that reduces the problem to a single-sample bit-scale optimization. This approach is analogous to the role of Stochastic Gradient Descent (SGD) or mini-batch techniques in classical deep learning, but is specifically tailored for our quantum-compatible optimization framework.
Here, we consider the following two-stage stochastic QCBO problem. The first-stage decision variables correspond to the neural network parameters, and are represented by the vector $\mathbf{x}\in\{0,1\}^{n}$. After $\mathbf{x}$ is chosen, one of $N$ samples is realized. For each sample $1\leq i\leq N$, a set of sample specific decisions $\mathbf{y}_{i}\in\{0,1\}^{m}$ can be made. The problem could be formulated as:
$\min\quad$ $\mathbf{x}^{\top}\mathbf{A}\mathbf{x}+\mathbf{a}^{\top}\mathbf{x}+\sum_{i=1}^{N}p_{i}\left(\mathbf{x}_{i}^{\top}\mathbf{B}_{i}\mathbf{y}_{i}+\mathbf{y}_{i}^{\top}\mathbf{C}_{i}^{\top}\mathbf{y}_{i}+\mathbf{c}_{i}^{\top}\mathbf{y}_{i}\right)$ (52)
s.t. $\mathbf{F}_{i}\mathbf{x}+\mathbf{G}_{i}\mathbf{y}_{i}=\mathbf{r}_{i},\quad\forall 1\leq i\leq N,$ (53)
$\mathbf{x}\in\{0,1\}^{n},\quad\mathbf{y}_{i}\in\{0,1\}^{m},\quad\forall 1\leq i\leq N$ (54)
Here, all variables are binary. The matrices $\mathbf{A}$, $\mathbf{B}_{i}$, $\mathbf{C}_{i}$, $\mathbf{F}_{i}$, $\mathbf{G}_{i}$ and vectors $\mathbf{a}$, $\mathbf{c}_{i}$, $\mathbf{r}_{i}$ are the data for the problem, and $p_{i}$ represents the probability of sample $i$.
As shown in Figure 3, the decomposed relaxation offers an alternative to the monolithic exact completely positive programming (CPP) formulation, aiming to mitigate the computational burden of handling a single large-scale cone constraint by distributing the problem across multiple sample-specific cones. This approach [38, 39] leverages a separable structure that facilitates scalability and enables the application of specialized quantum devices with limited qubits.
The objective function is designed to balance the first-stage decisions with sample dependent recourse actions, weighted by their respective probabilities. It is formulated as:
$\min\quad$ $\text{Tr}(\mathbf{A}\mathbf{X})+\mathbf{a}^{\top}\mathbf{x}+\sum_{i=1}^{N}p_{i}\left(\text{Tr}(\mathbf{B}_{i}\mathbf{Z}_{i})+\text{Tr}(\mathbf{C}_{i}\mathbf{Y}_{i})+\mathbf{c}_{i}^{\top}\mathbf{y}_{i}\right)$ (55)
Here, $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $\mathbf{B}_{i}\in\mathbb{R}^{m\times m}$ are symmetric matrices representing quadratic costs, $\mathbf{a}\in\mathbb{R}^{n}$ and $\mathbf{c}_{i}\in\mathbb{R}^{m}$ are linear cost vectors. The matrices $\mathbf{X}\in\mathbb{R}^{n\times n}$, $\mathbf{Y}_{i}\in\mathbb{R}^{m\times m}$ and $\mathbf{Z}_{i}\in\mathbb{R}^{m\times n}$ are lifted variables corresponding to $\mathbf{x}\mathbf{x}^{\top}$, $\mathbf{y}_{i}\mathbf{y}_{i}^{\top}$ and cross-term $\mathbf{y}_{i}\mathbf{x}^{\top}$, respectively.
The following constraints enforce the problem’s structural and conic properties, while decomposing the global cone into sample-specific components:
  • Linear Constraints: The original recourse constraints are preserved for each sample:
    $\mathbf{F}_{i}\mathbf{x}+\mathbf{G}_{i}\mathbf{y}_{i}=\mathbf{r}_{i},\quad\forall 1\leq i\leq N,$ (56)
    where $\mathbf{F}_{i}\in\mathbb{R}^{k\times n}$, $\mathbf{G}_{i}\in\mathbb{R}^{k\times m}$, and $\mathbf{r}_{i}\in\mathbb{R}^{k}$ are given data matrices and vectors.
  • Squared Linear Constraints: Each linear constraint is squared and expressed using the lifted matrix variables to capture second-order effects:
    $\begin{pmatrix}(\mathbf{F}_{i})_{k,:}&(\mathbf{G}_{i})_{k,:}\end{pmatrix}\begin{pmatrix}\mathbf{X}&\mathbf{Z}_{i}^{\top}\\ \mathbf{Z}_{i}&\mathbf{Y}_{i}\end{pmatrix}\begin{pmatrix}(\mathbf{F}_{i})_{k,:}^{\top}\\ (\mathbf{G}_{i})_{k,:}^{\top}\end{pmatrix}=r_{i,k}^{2},\quad\forall 1\leq i\leq N,$ (57)
    where $(\mathbf{F}_{i})_{k,:}$ and $(\mathbf{G}_{i})_{k,:}$ denote the $k$-th rows of $\mathbf{F}_{i}$ and $\mathbf{G}_{i}$, respectively, and $r_{i,k}$ denotes the $k$-th entry of $\mathbf{r}_{i}$.
  • Binary Constraints: The binary nature of the variables is enforced by linking the vector variables to the diagonals of the matrix variables:
    $x_{\ell}=\mathbf{X}_{\ell\ell},\quad y_{i,\ell}=\mathbf{Y}_{i,\ell\ell},\quad\forall 1\leq\ell\leq n,\forall 1\leq i\leq N,$ (58)
  • Decomposed Cone Constraints: A sample-specific completely positive cone constraint is imposed on the augmented matrix for each $i$:
    $\begin{pmatrix}1&\mathbf{x}^{\top}&\mathbf{y}_{i}^{\top}\\ \mathbf{x}&\mathbf{X}&\mathbf{Z}_{i}^{\top}\\ \mathbf{y}_{i}&\mathbf{Z}_{i}&\mathbf{Y}_{i}\end{pmatrix}\in\mathcal{C}_{1+n+m}^{*},\quad\forall 1\leq i\leq N,$ (59)
    where $\mathcal{C}_{1+n+m}^{*}$ is the cone of $(1+n+m)\times(1+n+m)$ completely positive matrices, ensuring the nonconvex quadratic constraints are satisfied.
Fact 2 (Superiority of the DLBO Relaxation over the Classical SDP Relaxations, [39])
The Decomposed Lower Bound (DLBO) formulation (56)-(59) provides a tighter relaxation compared to the standard sparse Semidefinite Programming (SDP) relaxation, i.e., for a minimization problem, the optimal value obtained from the DLBO provides a lower bound that is greater than or equal to that from the SDP relaxation:
$\text{val}(\text{SDP})\leq\text{val}(\text{DLBO}).$ (60)
When applying QCGD algorithm to the decomposed relaxation problem above, in each step we need to solve the following problem,
$\min\quad$ $\sum_{i=1}^{N}\text{Tr}(\mathbf{G}_{t}\mathbf{P}_{i})$ (61)
s.t. $\mathbf{P}_{i}=\begin{pmatrix}1&\mathbf{u}^{\top}&\mathbf{v}_{i}^{\top}\\ \mathbf{u}&\mathbf{U}&\mathbf{W}_{i}^{\top}\\ \mathbf{v}_{i}&\mathbf{W}_{i}&\mathbf{V}_{i}\end{pmatrix}\in\mathcal{C}^{*}_{1+n+m},\quad\forall 1\leq i\leq N$ (62)
The QCGD subproblem (61)-(62) requires minimizing a linear function over the Cartesian product of $N$ completely positive cones, which are coupled through the shared parameters $\mathbf{u}$. The coupling through the block of variables $(\mathbf{u},\mathbf{U})$ and the separability of the blocks $(\mathbf{v}_{i},\mathbf{V}_{i},\mathbf{W}_{i})$ makes this problem amenable to a Block Coordinate Descent (BCD) approach [51], as completely positive cone is a special case of semidefinite cone. We decompose the variables into $N+1$ blocks: a public block $(\mathbf{u},\mathbf{U})$ and $N$ private blocks $(\mathbf{v}_{i},\mathbf{W}_{i},\mathbf{V}i)$, one per sample. The BCD algorithm then alternates between two phases: (i) optimizing the public block while fixing all private blocks, and (ii) optimizing each private block (in parallel) while fixing the public block. Each subproblem involves linear minimization over $\mathcal{C}^{*}_{1+n+m}$, which can be reformulated and solved using Ising hardware, according to the transformation we will justify later.
Then, we begin by initializing $\mathbf{P}_{i}^{(0)}$ for all $i$ with feasible, completely positive matrices. At each iteration, we perform the following operations:
  • Per-Sample Recourse. In this step, we fix the public variables at their current values, $\mathbf{u}=\mathbf{u}^{(k)}$ and $\mathbf{U}=\mathbf{U}^{(k)}$. The overall objective function decouples into $N$ independent subproblems. For a sample-specific block ($k=i$), fix $\mathbf{u}$, $\mathbf{U}$, and all other $\mathbf{v}_{j}$, $\mathbf{W}_{j}$, $\mathbf{V}_{j}$ for $j\neq i$, and optimize over $\mathbf{v}_{i}$, $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$. The subproblem is defined as $\min_{\mathbf{v}_{i},\mathbf{W}_{i},\mathbf{V}_{i}}\text{Tr}(\mathbf{G}_{i}\mathbf{P}_{i})$, subject to $\mathbf{P}_{i}\in\mathcal{C}_{1+n+m}^{*}$. This isolates the contribution of sample $i$ to the overall objective. The solution gives us the updated recourse variables $\mathbf{v}_{i}^{(k+1)}$.
  • Update Neural Network Parameters. Next, we fix the private recourse variables at their new values, i.e., $\mathbf{v}_{i}=\mathbf{v}_{i}^{(k+1)}$, $\mathbf{W}_{i}=\mathbf{w}_{i}^{(k+1)}$, and $\mathbf{V}_{i}=\mathbf{V}_{i}^{(k+1)}$, and optimize for the public parameters $\mathbf{u}$. The subproblem becomes: $\min_{\mathbf{u},\mathbf{U}}\sum_{i=1}^{N}\text{Tr}(\mathbf{G}_{i}\mathbf{P}_{i})$, subject to $\mathbf{P}_{i}$ remaining in $\mathcal{C}_{1+n+m}^{*}$.
From Conic Optimization to QUBO Formulation.
The reduction of each block optimization step to a QUBO is performed as follows. Using the equivalence between separation and optimization [52], we determine if $\mathbf{P}_{i}\in\mathcal{C}_{1+n+m}^{*}$, otherwise, a separating hyperplane $\mathbf{Z}\in\mathcal{C}_{1+n+m}$ exists such that $\text{Tr}(\mathbf{Z}\mathbf{P}_{i})<0$. Thus, the separation oracle can be obtained by solving the problem as $\min_{\mathbf{Z}\in\mathcal{C}_{1+n+m}}\text{Tr}(\mathbf{Z}\mathbf{P}_{i}).$ The problem now becomes an optimization over the copositive cone $\mathcal{C}_{1+n+m}$. A matrix $\mathbf{Z}$ is copositive if $\mathbf{x}^{\top}\mathbf{Z}\mathbf{x}\geq 0$ for all $\mathbf{x}\geq 0$. Using the separation-optimization equivalence again, the problem could be reduced to determining whether $\mathbf{Z}$ is copositive, which can be checked by solving $\min_{\mathbf{y}\geq 0}\mathbf{y}^{\top}\mathbf{Z}\mathbf{y}$ using Ising hardware. If the minimum is non-negative, $\mathbf{Z}$ is copositive; otherwise, a separating vector $\mathbf{y}$ is obtained.
Quantum Progressive Hedging (QPH) Algorithm.
Following the separable CP formulations, the sample-$i$ contribution to the objective is
$f_{i}(\mathbf{x},\mathbf{X},\mathbf{W}^{(i)}):=\;p_{i}\Big(\text{Tr}(\mathbf{B}_{i}\mathbf{Z}_{i})+\text{Tr}(\mathbf{C}_{i}\mathbf{Y}_{i})+\mathbf{c}_{i}^{\top}\mathbf{y}_{i}\Big),\qquad\mathbf{W}^{(i)}:=(\mathbf{y}_{i},\mathbf{Y}_{i},\mathbf{Z}_{i}).$ (63)
For a fixed neural network parameters $\mathbf{x},\mathbf{X}$, we can define the feasible set of scenario $i$ as:
[…]
The partial minimization (recourse value) for scenario $i$ is
$\phi_{i}(\mathbf{x},\mathbf{X}):=\inf_{\mathbf{W}^{(i)}\in\mathcal{K}_{i}(\mathbf{x})}f_{i}(\mathbf{x},\mathbf{X},\mathbf{W}^{(i)}).$ (65)
Intuitively, $\phi_{i}(\mathbf{x},\mathbf{X})$ is the best achievable (relaxed) loss on sample $i$ given network parameters $\mathbf{x},\mathbf{X}$. Under the conditions that $\mathcal{K}_{i}(\mathbf{x},\mathbf{X})$ is a convex set whose graph is convex in $(\mathbf{x},\mathbf{X},\mathbf{W}^{(i)})$, and $f_{i}$ in (63) is linear, the partial minimization $\phi_{i}(\cdot)$ is convex. Also, the global objective of DLBO becomes
$\min_{\mathbf{x},\mathbf{X}}\quad\text{Tr}(\mathbf{A}\mathbf{X})+\mathbf{a}^{\top}\mathbf{x}\;+\;\sum_{i=1}^{N}\phi_{i}(\mathbf{x},\mathbf{X}).$ (66)
In the QPH algorithm, we introduce per-sample copies $\mathbf{x}^{(i)},\mathbf{X}^{(i)}$ of the neural network parameters and enforce $\mathbf{x}^{(i)}=\bar{\mathbf{x}}$, $\mathbf{X}^{(i)}=\bar{\mathbf{X}}$, where $\bar{\mathbf{x}}$ and $\bar{\mathbf{X}}$ are a global consensus. Equivalently, we can rewrite Eq.(66) as
$\min_{\{\mathbf{x}^{(i)},\mathbf{X}^{(i)},\mathbf{W}^{(i)}\}_{i=1}^{N},\ \bar{\mathbf{x}},\bar{\mathbf{X}}}$ $\text{Tr}(\mathbf{A}\mathbf{X}^{(i)})+\mathbf{a}^{\top}\bar{\mathbf{x}}\;+\;\sum_{i=1}^{N}f_{i}(\mathbf{x}^{(i)},\mathbf{X}^{(i)},\mathbf{W}^{(i)})$ (67)
s.t. $\mathbf{W}^{(i)}\in\mathcal{K}_{i}(\mathbf{x}^{(i)},\mathbf{X}^{(i)}),\quad i=1,\dots,N,$
$\mathbf{x}^{(i)}=\bar{\mathbf{x}},\mathbf{X}^{(i)}=\bar{\mathbf{X}},\quad i=1,\dots,N.$
Let $\bm{\lambda}^{(i)}_{1}\in\mathbb{R}^{n}$, $\bm{\lambda}^{(i)}_{2}\in\mathbb{R}^{n\times n}$ be dual multipliers for $\mathbf{x}^{(i)}=\bar{\mathbf{x}}$, $\mathbf{X}^{(i)}=\bar{\mathbf{X}}$ respectively, and fix the penalty as $\rho>0$. The augmented Lagrangian contribution of sample $i$ at iteration $t$ could be expressed as
$\mathcal{L}^{(i)}_{\rho}\big(\mathbf{x}^{(i)},\mathbf{X}^{(i)},\mathbf{W}^{(i)};\,\bar{\mathbf{x}}_{t},\bar{\mathbf{X}}_{t},\bm{\lambda}^{(i)}_{1,t},\bm{\lambda}^{(i)}_{2,t}\big)$
$:=$ $f_{i}(\mathbf{x}^{(i)},\mathbf{X}^{(i)},\mathbf{W}^{(i)})+(\bm{\lambda}^{(i)}_{1,t})^{\top}(\mathbf{x}^{(i)}-\bar{\mathbf{x}}_{t})+\text{Tr}((\bm{\lambda}^{(i)}_{2,t})^{\top}(\mathbf{X}^{(i)}-\bar{\mathbf{X}}_{t}))+\frac{\rho}{2}\,\big(\|\mathbf{x}^{(i)}-\bar{\mathbf{x}}_{t}\|_{2}^{2}+\|\mathbf{X}^{(i)}-\bar{\mathbf{X}}_{t}\|_{2}^{2}\big),$ (68)
with $\mathbf{W}^{(i)}\in\mathcal{K}_{i}(\mathbf{x}^{(i)},\mathbf{X}^{(i)})$. Then, the QPH performs the following updates:
(i) Subproblem for each sample (parallel over $i$).
For each sample $i$, it is required to solve
[…]
Each problem in Eq.(69) involves only one sample $i$ and its lifted CP block $\mathbf{P}^{(i)}$. Using the conic-to-QUBO reduction outlined previously, we map the equation to a QUBO instance and then repeatedly call a quantum Ising solver (such as a coherent Ising machine or a quantum annealer). Among the returned bit strings, we select the candidate with the lowest energy as the solution.
(ii) Consensus update.
After all samples are processed, we can update the global non-anticipative network as the probability-weighted average $(\bar{\mathbf{x}}_{t+1},\bar{\mathbf{X}}_{t+1})\;\leftarrow\;\sum_{i=1}^{N}p_{i}\,(\mathbf{x}^{(i)}_{t+1},\mathbf{X}^{(i)}_{t+1})$. This is to minimization of the term $\sum_{i}p_{i}\tfrac{\rho}{2}\big(\|\mathbf{x}^{(i)}-\bar{\mathbf{x}}_{t}\|_{2}^{2}+\|\mathbf{X}^{(i)}-\bar{\mathbf{X}}_{t}\|_{2}^{2}\big)$.
(iii) Dual (hedging) update.
Finally, we have
$\bm{\lambda}^{(i)}_{1,t+1}\;\leftarrow\;\bm{\lambda}^{(i)}_{1,t}+\rho\,(\mathbf{x}^{(i)}_{t+1}-\bar{\mathbf{x}}_{t+1}),\quad\bm{\lambda}^{(i)}_{2,t+1}\;\leftarrow\;\bm{\lambda}^{(i)}_{2,t}+\rho\,(\mathbf{X}^{(i)}_{t+1}-\bar{\mathbf{X}}_{t+1}),\qquad i=1,\dots,N.$ (70)
Intuitively, $\bm{\lambda}^{(i)}_{1}$ and $\bm{\lambda}^{(i)}_{2}$ penalize sample $i$ for deviating from the global consensus network.
The convergence of the QPH algorithm adheres to the standard theoretical framework of stochastic programs [53] and consensus ADMM theory [54]. The quadratic penalty term in the augmented Lagrangian formulation (Quantum Progressive Hedging (QPH) Algorithm.) renders each sample subproblem strongly convex in $\mathbf{x}^{(i)}$, thereby guaranteeing its well-posedness. The dual variable update (70) is designed to progressively enforce the non-anticipativity constraint. As established in the literature, ADMM achieves an $O(1/t)$ convergence rate for both primal feasibility violation and the suboptimality gap [54]. Under additional regularity assumptions, such as strong convexity and Lipschitz gradients, the consensus residual $\mathbf{x}^{(i)}_{t}-\bar{\mathbf{x}}_{t}||_{2}$ is proven to converge linearly [55].
Experimental validations
To verify the advantages of the training on neural networks, we undertake the experiment using coherent Ising machines for the classification of coat and sandal images from the Fashion MNIST dataset. The dataset consists of $60,000$ training images and $10,000$ test images, each representing a grayscale clothing item from $10$ classes, including T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. Each image is $28\times 28$ pixels, with pixel values ranging from 0 to $255$. Additional details of the experimental results are deferred to Section 6.1 of the supplementary.
Figure 4(a) illustrates the energy evolution during the optimization process using CIM to solve the neural network training directly. CIM identifies the optimal solution at the 88th iteration, approaching the optimal value in 0.185 ms, with the resulting model achieving a test accuracy of $94.95\%$.
To evaluate the performance of our quantum-based approach for training quantized neural networks, we present a series of comparative analyses. Figure 4(b) compares the test accuracy and running time of our model and quantum approach against previous classical approaches, including straight through estimator (STE) [56] and BinaryConnect [57]. Here, the proposed quantum-based method achieves a test accuracy of $94.95\%$ with a significantly reduced running time of $0.185$ ms, demonstrating superior efficiency and effectiveness.
Figure 4: (a) Energy evolution curve during the optimization of a quantized neural network using a coherent Ising machine (CIM); (b) Comparative analysis of test accuracy and running time for different training algorithms (STE [56], BinaryConnect [57]); (c) Performance comparison of memory demand, inference time, and accuracy before and after quantization and activation function approximation.
Further, Figure 4(c) examines the memory demand and inference complexity, contrasting a neural network with continuous weights and sigmoid activation functions against our quantized approach, which approximates the activation function with a simple piecewise linear function. Our method reduces memory cost to 2.75 bytes and inference time to 0.0762 seconds while maintaining a test accuracy of $94.95\%$, highlighting its resource efficiency. These results collectively underscore the advantages of leveraging quantum computing for training quantized neural networks, offering speed improvements and resource efficiency.
During inference with piecewise linear activation, the activation value at each neuron is determined in two steps: first by computing the pre-activation input, and then by identifying the interval within which this input falls. This interval-indexing step, or equivalently, locating the unique $\beta_{i}=1$, allows the model to evaluate the activation function through the corresponding piecewise linear segment. This process is computationally more efficient than evaluating a smooth nonlinear activation function (e.g., sigmoid or $\tanh$), which typically involves multiple floating-point operations such as exponentiation, division, and additional multiplications arising from polynomial or rational approximations used in hardware implementations.
In contrast, once the correct interval is determined, a piecewise linear activation requires only one interval lookup, one multiplication, and one addition. The time per activation evaluation can therefore be expressed as $T_{\mathrm{PWL}}=T_{\mathrm{lookup}}+T_{\mathrm{mul}}+T_{\mathrm{add}},$ which is lower than the cost of evaluating a smooth nonlinear activation, such as a sigmoid in the form $T_{\mathrm{sigmoid}}=T_{\exp}+T_{\mathrm{add}}+T_{\mathrm{div}}$. Consequently, spline-based piecewise linear activations not only facilitate a QCBO formulation during training but also offer a tangible acceleration during inference.
Table 2 summarizes the performance of post-training quantization (PTQ) algorithms under different activation functions and bit widths. As expected, reducing the bit precision generally leads to a degradation in accuracy.
Algorithm Activation FP Model 4-bit 3-bit 2-bit 1-bit
RTN [58] ReLU 98.8% 97.73% 97.73% 73.35% 50.00%
LeakyReLU 98.8% 98.70% 98.62% 76.77% 50.00%
Sigmoid 98.8% 98.76% 98.34% 55.84% 50.00%
AdaRound [9] ReLU 98.8% 97.73% 97.76% 69.87% 50.00%
LeakyReLU 98.8% 98.68% 98.78% 71.86% 50.00%
Sigmoid 98.8% 98.77% 98.61% 66.60% 50.00%
DFQ [59] ReLU 98.8% 96.62% 96.07% 84.99% 64.10%
LeakyReLU 98.8% 98.08% 97.80% 88.50% 61.40%
Sigmoid 98.8% 98.75% 96.44% 60.80% 50.00%
GPTQ [60] ReLU 98.8% 97.65% 97.50% 92.84% 50.00%
LeakyReLU 98.8% 98.80% 98.11% 91.82% 50.01%
Sigmoid 98.8% 98.80% 98.64% 91.21% 50.00%
ZeroQuant [61] ReLU 98.8% 97.79% 97.24% 93.59% 82.50%
LeakyReLU 98.8% 98.80% 98.10% 96.23% 91.43%
Sigmoid 98.8% 98.80% 97.28% 91.97% 91.45%
SubsetQ [62] ReLU 98.8% 97.79% 97.78% 96.70% 88.90%
LeakyReLU 98.8% 98.80% 98.65% 95.94% 91.45%
Sigmoid 98.8% 98.80% 98.80% 94.98% 91.45%
Table 2: Classification accuracy (%) under different PTQ algorithms, activation functions, and bit widths.
Across all post-training quantization baselines evaluated, including RTN [58], AdaRound [9], DFQ [59], GPTQ [60], ZeroQuant [61], and SubsetQ [62], accuracy degrades as the bit-width approaches the extremely low-precision regime ($\leq$2 bits). Conventional 1-bit and 2-bit post-training quantization methods often lead to severe representational collapse, with most failing to surpass $90\%$ accuracy and many performing no better than random chance. Against this backdrop, our quantum-trained model maintains desirable decision boundaries even at 1.1‑bit precision, attaining an accuracy of $94.95\%$.
In the low bit width regime, SubsetQ delivers competitive performance. However, it suffers from several inherent limitations. Notably, its computational cost grows sharply with increasing bit-width, because the method must evaluate across a candidate quantization set whose size scales as $2^{b}$. This exponential dependence renders SubsetQ efficient only in low precision regime (e.g., 121–2 bits), and increasingly prohibitive at higher bit‑widths. In contrast, our quantum‑optimization approach does not rely on enumerating quantization candidates and scales more favorably with target precision, thereby enabling flexible training at arbitrary effective bit‑widths, including the 1.1‑bit configuration employed in our experiments.
Spline interpolation requires careful consideration of the number of intervals, as it directly influences qubit resource allocation and must be optimized for quantum hardware constraints. In our experimental setup, we partitioned the input range into four intervals, achieving an optimal trade-off between qubit efficiency and segmentation accuracy. The corresponding breakpoints, defined at [8,c,0,c,8][-8,-c,0,c,8], yielded the following results:
Value of $c$ 1 2 3 4 5 6 7
Classfication Accuracy 96.35% 98.7% 96.35% 94.95% 96.75% 91.35% 91.35%
Table 3: Accuracy values for different cases of $c$.
Note that QUBO problems may yield multiple optimal solutions, and the accuracy shown corresponds to one such solution.
Figure 5: The figure illustrates the convergence of the Quantum Conditional Gradient Descent (QCGD) algorithm using CIM (green and red) and Gurobi (cyan and purple) solvers over 321 iterations on a log-log scale. The left y-axis shows the objective residual, and the right y-axis shows the constraint residual, with lines representing the residuals for CIM (solid) and Gurobi (dashed).
Figure 5 illustrates the convergence behavior of the Quantum Conditional Gradient Descent algorithm when applied to the neural network training utilizing a hybrid quantum-classical approach with CIM and Gurobi solvers. The algorithm achieves convergence to the optimal solution within 321 iterations, with the objective residual and constraint residual presented on a logarithmic scale. To further validate the robustness and reproducibility of these results, we performed five independent experiments, achieving convergence to the optimal solution in 723, 338, 263, 432, and 427 iterations, respectively. These consistent outcomes across multiple trials confirm its reliability and reproducibility in practical applications. In Section 6.2 of the supplementary, we also investigate the impact of coefficient matrix rounding and robustness of hybrid approach.
Effectiveness of DLBO framework
After solving the DLBO relaxation, we obtain a fractional parameter vector $\bar{x}\in[0,1]^{n}$. To obtain a quantized network, in this experiment, we apply an independent randomized Bernoulli rounding, i.e., for each coordinate $j$, we draw $X_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j})$, and interpret $X_{j}\in\{0,1\}$ as the final value of the $j$-th parameter bit. In practice, we repeat this randomized rounding procedure $K$ times, resulting in $K$ candidate binary networks $X^{(1)},\dots,X^{(K)}$. For each candidate, we evaluate the recourse objective $F(X^{(k)})$, i.e., the sum of per–sample losses measuring the discrepancy between the predicted labels and the true labels), and we report the best value as $F_{\min}:=\min_{1\leq k\leq K}F\bigl(X^{(k)}\bigr)$. Theoretical analysis in Section 6.3 of the supplementary shows that this strategy yields a discrete solution whose objective value concentrates around the DLBO relaxation value $F_{\mathrm{DLBO}}$ and improves with increasing $K$ with high probability.
The predictive performance of the DLBO is summarized in Table 4. Across all three datasets, Fashion MNIST [63], Wine [64], and Digits [65], the discrete networks obtained via randomized rounding maintain high approximation ratios, indicating that the performance of the learned binary structures remains extremely close to the relaxed optimum. Correspondingly, the classification performance remains competitive, with accuracies in the range of $90.24\%–96.30\%$ and $90.15\%–94.94\%$ on the test and full datasets, respectively. Weighted precision, recall, and F1 scores exhibit similar behavior, all falling within a relatively high numerical range, demonstrating that the DLBO framework produces stable and well-behaved decision boundaries after rounding.
Table 4: Performance Comparison on Test and Full Datasets
Test Set Full Set
Fashion MNIST Wine Digits Fashion MNIST Wine Digits
Approximation Ratio $0.9414$ $1.0000$ $1.0000$ $0.9414$ $1.0000$ $1.0000$
Accuracy $0.9024$ $0.9167$ $0.9630$ $0.9015$ $0.9494$ $0.9460$
Weight Avg. precision $0.9025$ $0.9235$ $0.9641$ $0.9015$ $0.9518$ $0.9519$
Weight Avg. recall $0.9024$ $0.9167$ $0.9630$ $0.9015$ $0.9494$ $0.9460$
Weight Avg. F1 Score $0.9024$ $0.9168$ $0.9630$ $0.9015$ $0.9496$ $0.9458$
Methods
Copositive Programming Framework
The QCGD method is built on copositive programming (CP) [47], which is a mathematical optimization branch that has garnered significant attention for its applications in various optimization problems. In copositive programming, the objective is to optimize over the cone of copositive matrices, which are matrices for which the quadratic form yields nonnegative values over the nonnegative orthant. One key application of copositive programming is in transforming binary quadratic problems into completely positive problems. This transformation involves reformulating the standard quadratic problem as a CP program, enabling the optimization of a quadratic form over the Cartesian product of simplices.
Consider the following linear constrained quadratic binary programming problem,
$\min\mathbf{x}^{T}\mathbf{Q}\mathbf{x}$ (71)
$s.t.\mathbf{A}\mathbf{x}=\mathbf{b}.$ (72)
According to [49, 47], the quadratic objective and linear constraints of the quadratic binary optimization problem can be represented as linear functions of a copositive matrix, and the problems (71)-(72) can be converted to the completely positive program.
In practice, we often encounter constraints in quadratic binary programming problems that are not only linear but may also be quadratic or even of higher order. For example, in the inequality (11) of the previous model, quadratic terms appear, requiring us to reduce the order of the problem. Recognizing that the copositive programming (CP) framework is suitable for handling constrained problems and can tolerate additional constraints, we can perform variable transformations and add extra constraints to reduce the order of the problem.
Consider the following quadratic 0-11 programming problem with quadratic constraints,
$\min\;$ $\mathbf{x}^{T}\mathbf{Q}_{0}\mathbf{x}$ (73)
s.t. $\mathbf{A}\mathbf{x}=\mathbf{b}$ (74)
$\mathbf{x}^{T}\mathbf{Q}_{t}\mathbf{x}\leq c_{t},1\leq t\leq T.$ (75)
To linearize the quadratic constraint, we can use $x_{ij}\in\{0,1\}$ to represent the product of $x_{i}$ and $x_{j}$, by adding the following constraints: $x_{ij}\leq x_{i},x_{ij}\leq x_{j},x_{ij}\geq x_{i}+x_{j}-1.$ It can be seen that $x_{ij}=x_{i}\cdot x_{j}$ always holds under these constraints.
In addition, we have the following improved linearization procedure. Consider the term $z^{l}_{x,j,k}=w^{l}_{jk}a^{l-1}_{x,k}$ in QCBO model, which can be expressed as $z^{l}_{x,j,k}=\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\cdot w^{l}_{jk}\cdot\delta^{(\ell)}_{a^{l}_{x,k}}$. In our neural network training model, quadratic constraints arise from the $w^{l}_{jk}a^{l-1}_{x,k}$ terms. Note that the bilinear components involving $w^{l}_{jk}$ and $a^{l-1}_{x,k}$ share the common variable $w^{l}_{jk}$. To linearize these terms effectively, we can use the following constraints:
$\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\,\delta^{(\ell)}_{a^{l}_{x,k}}+(2^{\bar{\ell}+1}-2)w^{l}_{jk}-z^{l}_{x,j,k}$ $\leq 2^{\bar{\ell}+1}-2,\quad\forall x,j,k,$ (76)
$z^{l}_{x,j,k}$ $\leq\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\,\delta^{(\ell)}_{a^{l}_{x,k}},\quad\forall x,j,k,$ (77)
$z^{l}_{x,j,k}$ $\leq(2^{\bar{\ell}+1}-2)w^{l}_{jk},\quad\forall x,j,k.$ (78)
Note that here we treat $w^{\ell}_{jk}$ as binary variables; for integer-valued $w^{\ell}_{jk}$, the linearization process described above is still applicable. We only need to perform the same operation on the bits of the binary encoding of $w^{\ell}_{jk}$ and $a^{l-1}_{x,k}$.
Quantum-Conditional Gradient Descent and Lazy Implementation
QCGD algorithm runs for $T$ iterations, with the following steps at each iteration $t$:
  1. 1.
    Set the step size $\gamma_{t}=2/(\delta(t+1))$ and penalty parameter $\alpha_{t}=\alpha_{0}\sqrt{\delta t+1}$, where $\delta>0$ is a scaling factor and $\alpha_{0}=1$ by default.
  2. 2.
    Compute errors in satisfying equality constraints ($\mathbf{g}_{t}$) and inequality constraints ($\mathbf{g}_{t}^{\prime}$) based on the current solution matrix $\mathbf{V}_{t}$.
  3. 3.
    Form the gradient matrix $\mathbf{G}_{t}$ using the coefficient matrix in the copositive programming formulation and constraint errors above. To ensure convergence, solve the subproblem $\min_{\mathbf{w}\in\mathbb{Z}_{2}^{p}}\mathbf{w}^{\top}\mathbf{G}_{t}\mathbf{w}$ $m_{t}$ times on a quantum annealer, selecting the best solution as:
    $\mathbf{w}_{t}=\arg\min_{\mathbf{w}^{(k)},k=1,\ldots,m_{t}}\left\{\mathbf{w}^{(k)\top}\mathbf{G}_{t}\mathbf{w}^{(k)}\right\},$ (79)
    where $\mathbf{w}^{(k)}$ is the $k$-th solution. Set the update direction as $\mathbf{D}_{t}=\mathbf{w}_{t}\mathbf{w}_{t}^{\top}$.
  4. 4.
    Update the solution: $\mathbf{V}_{t+1}=(1-\eta_{t})\mathbf{V}_{t}+\eta_{t}\mathbf{D}_{t}$.
  5. 5.
    Update the dual variables $\mathbf{z}_{t}$ and $\mathbf{z}_{t}^{\prime}$ using the step size $\gamma$.
  6. 6.
    After $T$ iterations, extract a binary solution $\mathbf{x}$ from $\mathbf{V}$, either directly from its first column or by computing the top singular vector of a submatrix and projecting onto the feasible set (e.g., using the Hungarian algorithm for permutation constraints).
The practical implementation of the QCGD algorithm is challenged by the error rates associated with quantum computers, leading to inaccuracies in the computation results. This inaccuracy is particularly pronounced in the oracle, which plays a crucial role in the QCGD algorithm by providing direction in each iteration. The consideration of the inexact oracle in the context of QCGD is therefore crucial for several reasons. Firstly, due to the inherent limitations of quantum hardware, achieving perfect precision in quantum computations is challenging. The inexactness in the oracle reflects these limitations and underscores the need to develop algorithms that are robust to such imperfections. Secondly, the use of an inexact oracle also introduces a trade-off between computational accuracy and efficiency.
Definition 1 ($(\delta,\varepsilon)$-Inexact Oracle [66, 67])
Consider the problem
$\min_{\mathbf{V}}\ \mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}\mathbf{V}\right)$ (80)
over the cone of completely positive matrices. An oracle is called a $(\delta,\varepsilon)$-inexact oracle if at iteration $t$ it outputs a matrix $\mathbf{D}_{t}$ such that
$\mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}(\mathbf{D}_{t}-\mathbf{V}_{t})\right)\;\leq\;\delta\cdot\mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}(\mathbf{D}^{*}_{t}-\mathbf{V}_{t})\right)\;+\;\frac{\xi_{t}}{\sqrt{t}},$ (81)
where $\mathbf{D}^{*}_{t}$ denotes the exact optimal solution of the QUBO problem, $\delta\in(0,1]$ is a relative error parameter, and $\xi_{t}$ is a random additive error satisfying $\mathbb{E}[\xi_{t}]\leq\varepsilon$.
A quantum computer without any imperfections corresponds to the scenario where $\delta=1$ and $\varepsilon=0$. The following lemma indicates that convergence can still be achieved even if the single quantum computation acts as an inexact oracle. This theoretical framework ensures that the overall algorithm remains robust despite potential inaccuracies in the quantum computation. Theoretical proof follows from that in [49], and we include it in Section 5 of the supplementary for completeness.
Proposition 1
Under $(\delta,\varepsilon)$-Inexact oracle, the objective gap and infeasibility in the $T$-th iteration of the hybrid algorithm satisfies that
$\mathbb{E}[\text{Objective-gap}_{T}]=O\Big(\frac{(1+\varepsilon)}{\sqrt{\delta^{3/2}T}}\Big),\mathbb{E}[\text{Infeasibility}_{T}]=O\Big(\frac{(1+\varepsilon)}{\sqrt{\delta^{3/2}T}}\Big)$ (82)
Here $\text{Objective-gap}_{T}$ denotes the objective value of the solution obtained at the $T$-th iteration and the optimal objective value; $\text{Infeasibility}_{T}$ represents the distance from the solution obtained at the $T$-th iteration to the feasible region. In addition, the bound in (82) holds with probability at least $1-\frac{\max\mathrm{Var}[\xi_{s}]}{\varepsilon^{2}T}$.
To reduce unnecessary computation, we can adopt a lazy implementation strategy, where solving the QUBO problem is applied only when significant changes occur. This approach is motivated by the observation that, in later iterations, the magnitude of QUBO coefficient modifications often becomes negligible. We can determine whether solving the QUBO problem is necessary based on the spectral gap:
$\Delta_{t}=\min_{\mathbf{D}\neq\mathbf{D}^{*}_{t}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D})-\min_{\mathbf{D}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D})$ (83)
which is defined as the difference between the lowest and the second-lowest objective value of the QUBO problem. If the cumulative changes in the coefficients of the QUBO matrix fall below a certain threshold, i.e.,
[…]
Then, it implies that the perturbations in the solution landscape are insignificant. Consequently, the algorithm can bypass solving the latest QUBO problem. By leveraging this stability, potentially through a threshold-based criterion to detect negligible changes, computational overhead can be reduced, particularly in resource-constrained environments.
Conclusion
In conclusion, this paper introduces a QUBO-based and convex formulation framework that enables arbitrary activation and loss functions through spline interpolation, significantly expanding the applicability of quantum computing and the understanding of optimization landscape in neural network training. The theoretically derived upper bound on Ising spin requirements, along with the empirically validated convergence of our hybrid algorithm, even under randomized outputs and limited coefficient matrix precision, validates the robustness of our approach. We also establish scalable hybrid training algorithms that decomposes the global QCBO problem into tractable quantum subproblems while preserving the tightness of the solution. Comparisons with classical methods highlights the superior performance of the Ising machine in quantized neural network training, while the hybrid algorithm demonstrates consistent efficacy and stability. Meanwhile, our model offers promising potential for applications demanding trainable activation functions, paving the way for quantum-enhanced neural architectures. This work represents a pivotal advancement toward harnessing the full potential of quantum-accelerated machine learning in real-world implementations.
Supplementary Material
Appendix A Related Work
In addition to the aforementioned works, other research also explores the intersection of quantum computing and neural networks. Ref. [68] presents a novel training approach based on Adiabatic Quantum Computing (AQC), which utilizes principles of adiabatic evolution to address optimization challenges. The proposed universal AQC scheme is designed for implementation on gate-based quantum computers. A hybrid strategy introduced in Ref. [69] combines quantum and classical methods to accelerate the training of binary neural networks (BNNs). Its quantum component employs the HHL algorithm to solve linear systems of equations for linear regression within a single-layer BNN. The primary contribution of Ref. [70] lies in demonstrating the transfer of a trained artificial neural network to a quantum computing environment. By integrating deep-learned parameters and layer structures, the authors formulate a quadratic binary model suitable for quantum annealing that aligns with the behavior of a classical neural network.
Training quantized neural network.
Neural network quantization techniques can be broadly categorized into quantization-aware training (QAT) and post-training quantization (PTQ). In QAT, quantization effects are simulated during training so that the optimizer learns weights that are intrinsically robust to low-bit inference. Early QAT methods include BinaryConnect [57], which constrains weights to binary values during the forward pass while maintaining full-precision copies for gradient updates. A key ingredient enabling QAT is the straight-through estimator (STE) [56, 71], which approximates gradients through discrete operations by substituting their non-differentiable Jacobians with surrogate identities. While QAT generally achieves superior accuracy, it requires full retraining, extensive data access, and increased computational cost.
In contrast, PTQ methods quantize a pre-trained model without full retraining and often use only a small calibration set. A straightforward baseline is round-to-nearest (RTN), which simply scales weights and rounds each to the nearest fixed-point value [58]. More sophisticated PTQ methods optimize the quantization procedure itself. AdaRound, for example, learns optimal weight rounding offsets by formulating the rounding decision as a local loss minimization (approximated as a QUBO problem) and solving it with a continuous relaxation [9]. Other approaches introduce lightweight fine-tuning steps into PTQ. For instance, layer-wise differentiable calibration methods adjust weights by backpropagating on a small calibration set to minimize the error between each layer’s quantized outputs and its full-precision outputs [72]. The recent GPTQ algorithm pushes this further by using second-order information, which applies an approximate Hessian-based optimizer to choose quantized weight values that incur minimal loss increase [60], enabling PTQ of large transformers down to 343–4 bits with negligible accuracy degradation. Several PTQ techniques target transformer models specifically. ZeroQuant combines dynamic, per-token activation quantization to suppress outliers with fine-grained group-wise weight quantization to maintain model accuracy on language tasks [61]. Meanwhile, SubsetQ introduces a non-uniform quantizer that selects an optimal subset of quantization levels from a larger universal set, allowing flexible step sizes tuned to each layer’s distribution [62].
Collectively, many of these advances frame quantization as an optimization problem. Indeed, AdaRound’s method explicitly casts weight quantization as a QUBO instance [9]. These motivates our approach, which leverages quantum optimization to explore discrete quantization configurations. It’s important to note that these classical methods still operate within the standard non-convex training landscape. We aim to further reduce the quantization error beyond the reach of conventional techniques.
Convex formulation of neural network training.
The training of deep neural networks presents a formidable non-convex optimization challenge, limiting theoretical guarantees and interpretability [73]. A significant line of research has therefore sought to reformulate this problem within the framework of convex optimization, aiming to provide pathways to certified global optimality. For ReLU activation, Sahiner and Pilanci et al. [74, 73] employs convex duality to construct a semi-infinite program whose constraints are parameterized by the discrete activation sign patterns of the ReLU neurons. While exact, this formulation’s complexity scales exponentially with the rank of the data matrix, restricting its direct application to low-dimensional settings [74]. A second, distinct approach involves lifting the problem into a higher-dimensional matrix space, which circumvents the need for sign pattern enumeration [73].
Appendix B Error Bound of Piecewise Linear Approximation in FNN
Lemma 3
Consider an $L$-layer feedforward neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$, where $W_{l}\in\mathbb{R}^{m_{l}\times m_{l-1}}$ represents the weight matrix in the $l$-th layer, and is spectrally normalized [75]. $\sigma:\mathbb{R}\to\mathbb{R}$ is Lipschitz continuous with constant $L_{\sigma}\leq 1$. Let $\hat{f}$ denote the network with $\sigma$ replaced by a piecewise linear function $\hat{\sigma}$, satisfying:
$\sup_{z\in\mathbb{R}}|\sigma(z)-\hat{\sigma}(z)|\leq\epsilon_{\sigma},$ (85)
and Lipschitz constant $L_{\hat{\sigma}}\leq 1$. Assume inputs $x\in\mathcal{X}\subset\mathbb{R}^{d}$ satisfy $\|x\|_{2}\leq B$, and each layer has $m_{l}\leq m$ neurons. Then:
$\sup_{x\in\mathcal{X}}\|f(x)-\hat{f}(x)\|_{2}\leq\epsilon_{\sigma}\sqrt{m}L.$ (86)
If $\sigma$ is twice continuously differentiable with $\|\sigma^{\prime\prime}\|_{\infty}\leq M$, the number of segments $n$ in $\hat{\sigma}$ to achieve an error no more than $\epsilon$ scales as:
$n=O\left(\frac{1}{\sqrt{\epsilon}}\right).$ (87)
This result and analysis provide insight into the trade-offs between model efficiency and approximation accuracy, particularly relevant in the context of quantum-enhanced neural network training.
Proof: Define layer-wise activations:
$h_{0}=x,\quad h_{l}=\sigma(\mathbf{W}_{l}h_{l-1}+\mathbf{b}_{l}),\quad\hat{h}_{0}=x,\quad\hat{h}_{l}=\hat{\sigma}(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l}),\quad l=1,\dots,L-1.$ (88)
The output error is:
$\|f(x)-\hat{f}(x)\|_{2}=\|\mathbf{W}_{L}(h_{L-1}-\hat{h}_{L-1})\|_{2}\leq\|h_{L-1}-\hat{h}_{L-1}\|_{2},$ (89)
since $\|\mathbf{W}_{L}\|_{2}\leq 1$. Let $\delta_{l}=\|h_{l}-\hat{h}_{l}\|_{2}$. We bound $\delta_{l}$ recursively.
For layer $l$, consider:
$h_{l}-\hat{h}_{l}=\sigma(\mathbf{W}_{l}h_{l-1}+\mathbf{b}_{l})-\hat{\sigma}(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l}).$ (90)
Decompose as:
$h_{l}-\hat{h}_{l}=\left[\sigma(\mathbf{W}_{l}h_{l-1}+\mathbf{b}_{l})-\sigma(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})\right]+\left[\sigma(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})-\hat{\sigma}(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})\right].$ (91)
The first term, using $L_{\sigma}\leq 1$, is:
$\|\sigma(\mathbf{W}_{l}h_{l-1}+\mathbf{b}_{l})-\sigma(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})\|_{2}\leq L_{\sigma}\|\mathbf{W}_{l}h_{l-1}-\mathbf{W}_{l}\hat{h}_{l-1}\|_{2}\leq\|\mathbf{W}_{l}\|_{2}\delta_{l-1}\leq\delta_{l-1}.$ (92)
The second term, by the approximation error, satisfies:
$\|\sigma(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})-\hat{\sigma}(\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l})\|_{2}=\sqrt{\sum_{i=1}^{m_{l}}|\sigma(z_{i})-\hat{\sigma}(z_{i})|^{2}}\leq\sqrt{m_{l}}\epsilon_{\sigma}\leq\sqrt{m}\epsilon_{\sigma},$ (93)
where $z=\mathbf{W}_{l}\hat{h}_{l-1}+\mathbf{b}_{l}$. Thus:
$\delta_{l}\leq\delta_{l-1}+\sqrt{m}\epsilon_{\sigma}.$ (94)
Solving the recursion, we can obtain that $\delta_{l}\leq l\sqrt{m}\epsilon_{\sigma}$. For the output:
$\|f(x)-\hat{f}(x)\|_{2}\leq\delta_{L-1}\leq(L-1)\sqrt{m}\epsilon_{\sigma}\leq L\sqrt{m}\epsilon_{\sigma}.$ (95)
To determine $n$, assume $\sigma$ has $\|\sigma^{\prime\prime}\|_{\infty}\leq M$. Construct $\hat{\sigma}$ over [R,R][-R,R] with $n$ equal segments, each of width $h=\frac{2R}{n}$. The interpolation error is:
$|\sigma(z)-\hat{\sigma}(z)|\leq\frac{1}{8}h^{2}\|\sigma^{\prime\prime}\|_{\infty}=\frac{1}{8}\left(\frac{2R}{n}\right)^{2}M=\frac{R^{2}M}{2n^{2}}.$ (96)
Set $\frac{R^{2}M}{2n^{2}}\leq\epsilon$, we have
$n\geq\sqrt{\frac{R^{2}M}{2\epsilon}},\quad n=O\left(\frac{1}{\sqrt{\epsilon}}\right).$ (97)
$\Box$
Appendix C Spline Quantumization Protocol
Quantum computing faces the dual challenge of optimizing not only QUBO models but also highly nonlinear optimization problems that are prevalent in scientific and industrial applications. The significance of this challenge lies in the ubiquity of problems featuring intricate, non-linear relationships among variables in real-world scenarios. This fact underscores the urgency for quantum algorithms capable of navigating and optimizing complex, highly nonlinear landscapes.
Formally, we consider the following optimization problem:
$\min_{\mathbf{x}\in\{0,1\}^{n}}f(h(\mathbf{x})),$ (98)
where $f(\cdot)$ is a function of arbitrary form. This formulation expands the potential for quantum computing to address real-world optimization challenges. A powerful method for approximating highly nonlinear functions is spline interpolation. Splines are piecewise linear functions that provide a smooth and flexible fit to data points, thereby transforming complex optimization landscapes into more tractable forms suitable for quantum optimization algorithms. The spline fitting process involves the following steps:
  • Knot Selection. Choose a set of knots $\{M_{i}\}_{i=1}^{n}$ which are the points in the domain of $f(\cdot)$ where the piecewise polynomial segments will join. These knots should be chosen to adequately capture the variability of $f(\cdot)$ over the entire domain, as the placement of knots can significantly affect the accuracy of the spline approximation.
  • Spline Construction. Define linear functions $S_{i}(x)$ for each segment between consecutive knots.
  • Spline Approximation. Replace the original function $f(\cdot)$ with the spline function $S(\cdot)$. The spline function can be expressed as:
    $S(h(\mathbf{x}))=\sum_{i=1}^{n}S_{i}(h(\mathbf{x}))\cdot\beta_{i},$ (99)
    where $\beta_{i}\in\{0,1\}$ denotes whether $h(\mathbf{x})$ is in the $i$-th iterval.
C.1 Piecewise Constant Segment
An effective method for simplifying highly nonlinear functions is piecewise constant fitting. This approach approximates the function $f(\cdot)$ using constant segments, which can reduce the complexity of the optimization problem and facilitate the translation into a QUBO model.
$\min$ $\sum_{i=0}^{n}{{\beta_{i}}\cdot S(M_{i})}$ (100)
$s.t$ $\sum_{i=1}^{n}{\beta_{i}}\cdot M_{i-1}\leq h(\mathbf{x})\leq\sum_{i=1}^{n}{\beta_{i}}\cdot M_{i}$ (101)
$\sum_{i=1}^{n}{\beta_{i}}=1,\beta_{i}\in\{0,1\}$ (102)
Dealing with the inequality constraint.
We have now derived a QUBO model capable of approximately solving optimization problems with arbitrary objective functions in (C). However, it is important to note that in representing $\beta_{i}$, we introduced two inequality constraints. Here, we introduce the following two approaches:
  • Adding a penalty term $(\sum_{i=1}^{n}{\beta_{i}}M_{i-1}+s-h(\mathbf{x}))^{2}$, where $s\in[0,\Delta M]$, assuming all the intervals are of length $\Delta M$.
  • For $s\in[-\frac{1}{2},\frac{1}{2}]$, adding a penalty term
    $\sum_{i=1}^{n-1}{\beta_{i}\cdot\Big(\frac{2h(\mathbf{x})-(M_{i-1}+M_{i})}{2(M_{i}-M_{i-1})}+s\Big)^{2}}+\Big(1-\sum_{i=1}^{n-1}\beta_{i}\Big)\Big(\frac{2h(\mathbf{x})-(M_{n-1}+M_{n})}{2(M_{n}-M_{n-1})}+s\Big)^{2}.$ (103)
    It is important to note that the higher-order terms cancel out when the expression is expanded.
We have the following theorem that elucidates the correctness of the second approach.
Theorem 3
Let $\mathbf{\beta}^{*}$ be the optimal solution vector for the following optimization problem,
$\min_{\mathbf{\beta}}\Big\{\sum_{i=1}^{n}{\beta_{i}\cdot\Big(\frac{2h(\mathbf{x})-(M_{i-1}+M_{i})}{2(M_{i}-M_{i-1})}\Big)^{2}}\Big|\sum_{i=1}^{n}\beta_{i}=1\Big\}.$ (104)
Then $\mathbf{\beta}^{*}_{i}=1$ only at index $i=i_{\mathbf{x}}$, where $M_{i_{\mathbf{x}}-1}\leq h(\mathbf{x})\leq M_{i_{\mathbf{x}}}$, while all other $\mathbf{\beta}^{*}_{i}$ are equal to 0.
Proof: To establish this conclusion, it suffices to observe the following: For $i=i_{\mathbf{x}}$, $|2h(\mathbf{x})-(M_{i_{\mathbf{x}}-1}+M_{i_{\mathbf{x}}})|\leq M_{i_{\mathbf{x}}}-M_{i_{\mathbf{x}}-1}$ and hence $(\frac{2h(\mathbf{x})-(M_{i_{\mathbf{x}}-1}+M_{i_{\mathbf{x}}})}{2(M_{i_{\mathbf{x}}}-M_{i_{\mathbf{x}}-1})})^{2}\leq\frac{1}{4}$. And for $i\neq i_{\mathbf{x}}$, we have $(\frac{2h(\mathbf{x})-(M_{i-1}+M_{i})}{2(M_{i}-M_{i-1})})^{2}>\frac{1}{4}$. $\Box$
Theorem 3 establishes that the penalty term in the new model effectively guides the values of $\beta_{i}$ to accurately represent the interval in which $h(\mathbf{x})$ resides. In fact, this new penalty term represents the squared weighted sum of distances from $h(\mathbf{x})$ to the midpoints of each interval. Consequently, the minimization of this penalty term incentivizes $\beta_{i}$ to take values that precisely indicate the correct interval for $h(\mathbf{x})$.
Building upon the versatility of spline approximation in simplifying complex optimization landscapes, we use the following sign function as an example to illustrate the efficiency of our approach.
Example 1 (Sign function)
$\text{sign}(x)$ (11 if $x\geq 0$ else 1-1) can be represented as
$\mathrm{sign}(x)=\min_{\beta\in\{0,1\}}\Big\{\Big(2\beta-1\Big)+\lambda\cdot\Big[\beta\cdot\Big(\frac{2x-M}{2M}+s\Big)^{2}+(1-\beta)\cdot(\frac{2x+M}{2M}+s)^{2}\Big]\Big\},$ (105)
where $M=\sup|x|$ and $s\in[-\frac{1}{2},\frac{1}{2}]$ is a slack variable.
Compared with the model in ref.[76], we do not need to introduce the intermediate variable. Additionally, when representing the sign activation function, we do not need to introduce auxiliary variables for order reduction [76]. These optimizations can save $O(WDN\log W)$ bits, where $W,D,N$ denote the network width, network depth, and dataset size, respectively. Indeed, we can reduce the coefficients of the highest-order terms in the model’s qubit count by $50\%$.
Appendix D Details of Convex Formulation
We consider a fixed dataset $\mathcal{S}=\{(\mathbf{x}^{(s)},y^{(s)})\}_{s=1}^{N}$ and an $L$-layer feedforward network with widths $m_{0},\dots,m_{L}$.
McCormick linear inequalities.
We now write the four inequalities in the global form $\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}\mathbf{u}\leq\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}},$ with explicit block matrices as following.
For fixed $(s,\ell)$, we define
$\bm{\delta}^{(\ell)}_{W}\in\{0,1\}^{m_{\ell}m_{\ell-1}R_{\ell}},\quad\mathbf{a}^{(s,\ell-1)}\in\mathbb{R}^{m_{\ell-1}},\quad\mathbf{v}^{(s,\ell)}\in\mathbb{R}^{m_{\ell}m_{\ell-1}R_{\ell}},$ (106)
where $\mathbf{v}^{(s,\ell)}$ stacks $v^{(s,\ell)}_{jk,r}$ over $(j,k,r)$ in a consistent order, and $\bm{\delta}^{(\ell)}_{W}$ stacks the matching bits $\delta^{(\ell)}_{jk,r}$ in the same order. Define the Kronecker matrices
$\underline{\mathbf{A}}^{(s,\ell-1)}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\mathrm{Diag}\!\big(\underline{\mathbf{a}}^{(s,\ell-1)}\big)\,\otimes\,\mathbf{I}_{R_{\ell}},\quad\overline{\mathbf{A}}^{(s,\ell-1)}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\mathrm{Diag}\!\big(\overline{\mathbf{a}}^{(s,\ell-1)}\big)\,\otimes\,\mathbf{I}_{R_{\ell}},$ (107)
and
$\mathbf{P}_{a}:=\mathbf{I}_{m_{\ell}}\,\otimes\,\big(\mathbf{1}_{R_{\ell}}^{\top}\otimes\mathbf{I}_{m_{\ell-1}}\big)\ \in\ \mathbb{R}^{(m_{\ell}m_{\ell-1}R_{\ell})\times m_{\ell-1}}.$ (108)
For each row indexed by $(j,k,r)$, $\mathbf{P}_{a}$ selects $a^{(s,\ell-1)}_{k}$ and replicates it across bits.
Using these blocks, the four inequalities can be written compactly as one stacked system
$\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}\begin{bmatrix}\bm{\delta}^{(\ell)}_{W}\\ \mathbf{a}^{(s,\ell-1)}\\ \mathbf{v}^{(s,\ell)}\end{bmatrix}\;\leq\;\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}},$ (109)
where
$\mathbf{N}^{(s,\ell)}_{\mathrm{Mc}}=\begin{bmatrix}-\overline{\mathbf{A}}^{(s,\ell-1)}&\mathbf{0}&\ \ \mathbf{I}\\ \underline{\mathbf{A}}^{(s,\ell-1)}&\mathbf{0}&-\mathbf{I}\\ -\underline{\mathbf{A}}^{(s,\ell-1)}&-\mathbf{P}_{a}&\ \ \mathbf{I}\\ \ \ \overline{\mathbf{A}}^{(s,\ell-1)}&\ \ \mathbf{P}_{a}&-\mathbf{I}\end{bmatrix},\qquad\mathbf{I}:=\mathbf{I}_{m_{\ell}m_{\ell-1}R_{\ell}}.$ (110)
The right-hand side vector $\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}}$ stacks the corresponding bound terms:
$\mathbf{n}^{(s,\ell)}_{\mathrm{Mc}}=\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ -\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}\\ \ \ \overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}\end{bmatrix}$, (111)
where
$\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}:=\Big(\mathbf{I}_{m_{\ell}}\otimes(\mathbf{1}_{R_{\ell}}\otimes\mathbf{I}_{m_{\ell-1}})\Big)\,\underline{\mathbf{a}}^{(s,\ell-1)},\qquad\overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}:=\Big(\mathbf{I}_{m_{\ell}}\otimes(\mathbf{1}_{R_{\ell}}\otimes\mathbf{I}_{m_{\ell-1}})\Big)\,\overline{\mathbf{a}}^{(s,\ell-1)}.$ (112)
Each block row corresponds exactly to one of the four scalar McCormick envelopes,
  • Row block 1 enforces $v^{(s,\ell)}_{jk,r}-\overline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq 0$.
  • Row block 2 enforces $v^{(s,\ell)}_{jk,r}+\underline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq 0$.
  • Row block 3 enforces $v^{(s,\ell)}_{jk,r}-a^{(s,\ell-1)}_{k}-\underline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq-\underline{a}^{(s,\ell-1)}_{k}$, which appears as $\underline{\mathbf{A}}^{(s,\ell-1)},\ -\mathbf{P}_{a},\ \mathbf{I}]\leq-\underline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}$.
  • Row block 4 enforces $v^{(s,\ell)}_{jk,r}+a^{(s,\ell-1)}_{k}+\overline{a}^{(s,\ell-1)}_{k}\delta^{(\ell)}_{jk,r}\leq\overline{a}^{(s,\ell-1)}_{k}$, which appears as $\overline{\mathbf{A}}^{(s,\ell-1)},\ \mathbf{P}_{a},\ -\mathbf{I}]\leq\overline{\mathbf{a}}^{(s,\ell-1)}_{\mathrm{rep}}$.
Reconstructing the affine layer.
Let $\mathbf{C}^{(\ell)}_{\mathrm{bit}}\in\mathbb{R}^{m_{\ell}m_{\ell-1}\times m_{\ell}m_{\ell-1}R_{\ell}}$ place the bit coefficient $c^{(\ell)}_{jk,r}$ at position $(j,k,r)$, and define
$\mathbf{u}^{(s,\ell)}_{\!\mathrm{bil}}\ =\ \mathbf{C}^{(\ell)}_{\mathrm{bit}}\mathbf{v}^{(s,\ell)}+\big(\mathbf{I}_{m_{\ell}}\otimes\mathrm{Diag}(\mathbf{a}^{(s,\ell-1)})\big)\mathbf{d}^{(\ell)}_{W},\qquad\mathbf{z}^{(s,\ell)}\ =\ \underbrace{\big(\mathbf{I}_{m_{\ell}}\otimes\mathbf{1}_{m_{\ell-1}}^{\top}\big)}_{\mathbf{S}^{(\ell)}_{\mathrm{sum}}}\mathbf{u}^{(s,\ell)}_{\!\mathrm{bil}}+\mathbf{b}^{(\ell)}.$ (113)
All relations in (113) are linear. The corresponding rows are assembled into a block-sparse matrix $\mathbf{M}_{\mathrm{aff}}$ acting on $\mathbf{u}$.
Activation and output selection.
Fix a grid $M_{0}<\cdots<M_{n}$, we define $\mathbf{M}_{L}=[M_{0},\dots,M_{n-1}]$, $\mathbf{M}_{U}=[M_{1},\dots,M_{n}]$, and $\mathbf{t}=[(\sigma(M_{i-1})+\sigma(M_{i}))/2]_{i}$. For each neuron $(s,\ell,j)$, we have
  • One-hot normalization:
    [ $\underbrace{\big[\ \mathbf{1}^{\top}\ \big]}_{\mathbf{m}^{\top}_{\text{1hot}}}\,\bm{\beta}^{(s,\ell)}_{z,j}\;=\;\underbrace{1}_{m_{\text{1hot}}}.$ (114)
    This row contributes to the global equality system $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q}$ by selecting the block $\bm{\beta}^{(s,\ell)}_{z,j}$ in $\mathbf{u}$ with coefficient vector $\mathbf{1}^{\top}$, and zero elsewhere.
  • Activation value tying:
    $\underbrace{\big[\ -(\mathbf{t}^{\top})\ \ \ 1\ \big]}_{\mathbf{m}^{\top}_{\text{act}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ a^{(s,\ell)}_{j}\end{bmatrix}\;=\;0.$ (115)
    Equivalently $a^{(s,\ell)}_{j}-\mathbf{t}^{\top}\bm{\beta}^{(s,\ell)}_{z,j}=0$. This also enters $\mathbf{M}_{q}\mathbf{u}=\mathbf{m}_{q}$ with right-hand side 0.
  • Lower and upper interval bound:
    $\underbrace{\big[\ -\mathbf{M}_{L}\ \ \ 1\ \big]}_{\mathbf{n}^{\top}_{\text{low}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ z^{(s,\ell)}_{j}\end{bmatrix}\;\geq\;0,\quad\text{i.e.}\quad z^{(s,\ell)}_{j}-\mathbf{M}_{L}\bm{\beta}^{(s,\ell)}_{z,j}\geq 0.$ (116)
    $\underbrace{\big[\ \mathbf{M}_{U}\ \ \ -1\ \big]}_{\mathbf{n}^{\top}_{\text{up}}}\,\begin{bmatrix}\bm{\beta}^{(s,\ell)}_{z,j}\\ z^{(s,\ell)}_{j}\end{bmatrix}\;\geq\;0,\quad\text{i.e.}\quad\mathbf{M}_{U}\bm{\beta}^{(s,\ell)}_{z,j}-z^{(s,\ell)}_{j}\geq 0.$ (117)
    These two rows together form the neuron-local block of the global inequality system $\mathbf{N}_{r}\mathbf{u}\leq\mathbf{n}_{r}$. Each such inequality will later receive a nonnegative slack variable $s_{r}$ so that it can be rewritten as an equality $\mathbf{N}_{r}\mathbf{u}+s_{r}=\mathbf{n}_{r}$ before CP lifting.
For the network output $a^{(s,L)}$, we introduce another one-hot vector $\bm{\beta}^{(s)}_{a}$ and apply the same construction, and the corresponding rows contribute additional blocks in the convex formulation in the same way.
Appendix E Proof of Convergence of QCGD with Random Quantum Oracle
Proof: Similar to ref.[49], the algorithm utilizes an augmented Lagrangian function, defined as
$\mathcal{Q}_{\alpha}(\mathbf{V},\mathbf{z})=\mathrm{Tr}(\mathbf{C}\mathbf{V})+\mathbf{z}^{\top}(\mathcal{L}\mathbf{V}-\mathbf{v})+\frac{\alpha}{2}\|\mathcal{L}\mathbf{V}-\mathbf{v}\|^{2}\quad\text{for}\penalty 10000\ \mathbf{V}\in\Delta^{p},$ (118)
where $\mathbf{V}$ represents the primal variable, $\mathbf{z}$ is the dual variable, $\alpha$ is the penalty parameter, $\mathbf{C}$ is the cost matrix, $\mathcal{L}$ denotes the matrix for linear constraints, and $\mathbf{v}$ is the vector of constraint values.
The QCGD algorithm begins by initializing $\alpha$, $\mathbf{V}$, and $\mathbf{z}$, and then proceeds iteratively through alternating primal and dual updates. In the primal step, the algorithm computes the gradient of the augmented Lagrangian with respect to $\mathbf{V}$ and identifies a direction that minimizes the linearized loss, which involves solving a standard QUBO problem to determine the update direction. The primal variable $\mathbf{V}$ is subsequently adjusted by moving along this direction. In the dual step, the gradient of the augmented Lagrangian with respect to $\mathbf{z}$ is computed, and $\mathbf{z}$ is updated using gradient ascent. Throughout the iterations, the penalty parameter $\alpha$ is increased as $\alpha_{i}=\alpha_{0}\sqrt{\delta i+1}$ to ensure that $\mathbf{V}$ converges to a feasible solution that satisfies the constraints.
We can obtain the following inequalities:
  • The smoothness of $\mathcal{Q}_{\alpha_{t}}$:
    $\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t+1},\mathbf{z}_{t})\leq\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t},\mathbf{z}_{t})+\gamma_{t}\cdot\delta\cdot\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}(\mathbf{V}_{\star}-\mathbf{V}_{t}))+O\Big(\alpha_{t}\gamma_{t}^{2}+\frac{\xi_{t}\gamma_{t}}{\sqrt{t}}\Big).$ (119)
  • Definition of $\mathbf{Q}^{(t)}_{\text{QUBO}}$:
    $\gamma_{t}\cdot\delta\cdot\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}(\mathbf{V}_{\star}-\mathbf{V}_{t}))\leq\gamma_{t}\cdot\delta\cdot\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})-\gamma_{t}\cdot\delta\cdot\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t},\mathbf{z}_{t})-\gamma_{t}\cdot\delta\cdot\frac{\alpha_{t}}{2}\|\mathcal{L}\mathbf{V}_{t}-\mathbf{v}\|^{2}.$ (120)
Combining (119) and (120), we have
$\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t+1},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})$
$\leq$ $(1-\gamma_{t}\cdot\delta)\Big(\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})\Big)-\gamma_{t}\cdot\delta\cdot\frac{\alpha_{t}}{2}\|\mathcal{L}\mathbf{V}_{t}-\mathbf{v}\|^{2}+O\Big(\alpha_{t}\gamma_{t}^{2}+\frac{\xi_{t}\gamma_{t}}{\sqrt{t}}\Big)$ (121)
Note that
$\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t},\mathbf{z}_{t})=\mathcal{Q}_{\alpha_{t-1}}(\mathbf{V}_{t},\mathbf{z}_{t})+\frac{\alpha_{t}-\alpha_{t-1}}{2}\|\mathcal{L}\mathbf{V}_{t}-\mathbf{v}\|^{2}$ (122)
Hence
$\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t+1},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})$
$\leq$ $(1-\gamma_{t}\cdot\delta)\Big(\mathcal{Q}_{\alpha_{t-1}}(\mathbf{V}_{t},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})\Big)+((1-\gamma_{t}\cdot\delta)(\alpha_{t}-\alpha_{t-1})/2-\gamma_{t}\cdot\delta\cdot\frac{\alpha_{t}}{2})\|\mathcal{L}\mathbf{V}_{t}-\mathbf{v}\|^{2}+O\Big(\alpha_{t}\gamma_{t}^{2}+\frac{\xi_{t}\gamma_{t}}{\sqrt{t}}\Big)$ (123)
By choosing $\gamma_{i}=\frac{2}{\delta\cdot(i+1)}$ and $\alpha_{i}=\alpha_{0}\sqrt{\delta i+1}$, we claim that
$(1-\gamma_{t}\cdot\delta)(\alpha_{t}-\alpha_{t-1})-\frac{\gamma_{t}\alpha_{t}}{2}\cdot\delta\leq 0.$ (124)
To verify this, first observe that
$1-\gamma_{t}\delta=\frac{t-1}{t+1},\quad\alpha_{t}-\alpha_{t-1}=\alpha_{0}\frac{\delta}{\sqrt{\delta t+1}+\sqrt{\delta(t-1)+1}}.$ (125)
The LHS becomes:
$\frac{\alpha_{0}}{t+1}\left(\frac{(t-1)\delta}{\sqrt{\delta t+1}+\sqrt{\delta(t-1)+1}}-\sqrt{\delta t+1}\right)$ (126)
$=$ $\frac{\alpha_{0}}{t+1}\left(\frac{-\delta-1-\sqrt{(\delta t+1)(\delta(t-1)+1)}}{\sqrt{\delta t+1}+\sqrt{\delta(t-1)+1}}\right)\leq 0$ (127)
Similar as [49], we can obtain the following bound:
$\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{t+1},\mathbf{z}_{t+1})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})\leq(1-\gamma_{t}\cdot\delta)\Big(\mathcal{Q}_{\alpha_{t-1}}(\mathbf{V}_{t},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})\Big)+O\Big(\alpha_{t}\gamma_{t}^{2}+\frac{\xi_{t}\gamma_{t}}{\sqrt{t}}\Big).$ (128)
Define the error $e_{t}=\mathcal{Q}_{\alpha_{t-1}}(\mathbf{V}_{t},\mathbf{z}_{t})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})$. The recursion becomes:
$e_{t+1}\leq(1-\gamma_{t}\delta)e_{t}+O\left(\alpha_{t}\gamma_{t}^{2}+\frac{\xi_{t}\gamma_{t}}{\sqrt{t}}\right).$ (129)
Note that the terms:
  • $1-\gamma_{t}\delta=1-\frac{2}{t+1}=\frac{t-1}{t+1}$.
  • $\alpha_{t}\gamma_{t}^{2}=\alpha_{0}\sqrt{\delta t+1}\cdot\frac{4}{\delta^{2}(t+1)^{2}}\sim\frac{4\alpha_{0}}{\delta^{3/2}}t^{-3/2}$.
  • $\frac{\varepsilon\gamma_{t}}{\sqrt{t}}=\varepsilon\cdot\frac{2}{\delta(t+1)}\cdot\frac{1}{\sqrt{t}}\sim\frac{2\varepsilon}{\delta}t^{-3/2}$.
  • Total perturbation: $O\left(\left(\frac{4\alpha_{0}}{\delta^{3/2}}+\frac{2\varepsilon}{\delta}\right)t^{-3/2}\right)$.
Iterate the inequality:
$e_{t+1}\leq$ $\prod_{s=1}^{t}(1-\gamma_{s}\delta)e_{1}+\sum_{s=1}^{t}\left[\prod_{k=s+1}^{t}(1-\gamma_{k}\delta)\right]O\left(\alpha_{s}\gamma_{s}^{2}+\frac{\xi_{s}\gamma_{s}}{\sqrt{s}}\right)$ (130)
$=$ $O\left(\frac{\frac{4\alpha_{0}}{\delta^{3/2}}}{t^{1/2}}\right)+O\Big(\frac{2}{\delta t^{2}}\cdot\underbrace{\sum_{s=1}^{t}\xi_{s}s^{1/2}}_{S(t)}\Big).$ (131)
For the additive error term,
$\mathbb{E}[S(t)]\leq\sum_{s=1}^{t}\varepsilon s^{1/2}=O(\varepsilon t^{3/2}).$ (132)
According to Chebyshev’s inequality,
$\mathbb{P}(S(t)\geq\mathbb{E}[S(t)]+\varepsilon t^{3/2})\leq\frac{\mathrm{Var}[S(t)]}{\varepsilon^{2}t^{3}}\leq\frac{\max\mathrm{Var}[\xi_{s}]}{\varepsilon^{2}t}.$ (133)
Therefore
$\mathbb{E}[e_{t}]=O\Big(\frac{\frac{4\alpha_{0}}{\delta^{3/2}}+\frac{2\varepsilon}{\delta}}{t^{1/2}}\Big)$ (134)
and
$\mathbb{P}\Big[e_{t}=O\Big(\frac{\frac{4\alpha_{0}}{\delta^{3/2}}+\frac{2\varepsilon}{\delta}}{t^{1/2}}\Big)\Big]\geq 1-\frac{\max\mathrm{Var}[\xi_{s}]}{\varepsilon^{2}t}.$ (135)
Consequently,
$\mathbb{P}[\lim_{t\rightarrow\infty}e_{t}=0]=1.$
Similar to the proof in [49],
$\mathcal{Q}_{\alpha_{t}}(\mathbf{V}_{T+1},\mathbf{z}_{T+1})\geq\mathrm{Tr}(\mathbf{C}\mathbf{V}_{T+1})-\frac{D^{2}}{2\alpha_{t}},$ (136)
We then have
$\text{Objective-gap}_{T}=\mathrm{Tr}(\mathbf{C}\mathbf{V}_{T+1})-\mathrm{Tr}(\mathbf{C}\mathbf{V}_{\star})=O\Big(\frac{(1+\varepsilon)}{\delta^{3/2}\sqrt{T}}\Big)$ (137)
For the infeasibility bound, the proof is similar as [49], we can derive the following bound:
$\text{Infeasibility}_{T}=\|\mathcal{L}\mathbf{V}_{T+1}-\mathbf{v}\|=O\Big(\frac{(1+\varepsilon)}{\delta^{3/2}\sqrt{T}}\Big).$ (138)
$\Box$
Appendix F Additional Experimental Results
Figure 6: (a) Circular heatmaps depict the evolution of the QUBO matrix during the QCGD algorithm for different significant digit precisions (1 to 6 digits) across selected iterations (300, 600, 900, 1200, 1500, 1800, 2100, 2400, 2700, and 3000). The color intensity represents matrix element values, with the color bar indicating the range from 0.0 (purple) to 1.0 (yellow). Each row corresponds to a specific precision level, illustrating the iterative trajectory and convergence behavior. (b) and (c): The two subplots show the smoothed absolute differences $|\mathrm{obj}_{k}-\mathrm{obj}_{10}|$ (left) and $|\mathrm{residual}_{k}-\mathrm{residual}_{10}|$ (right) relative to the 10-digit precision baseline across iterations for precision levels of 1 to 9 digits. The y-axis is on a log scale, with dashed lines representing each precision level.
F.1 Additional Details of the Fashion MNIST Binary Classification Results
The training set consists of $12000$ images ($6000$ coat and $6,000$ sandal images), and the test set consists of $2000$ images ($1000$ coat and $1000$ sandal images). Each image is divided into $3$ columns vertically, resulting in two $28\times 9$ grids and one $28\times 10$ grid. We then calculate the number of zero pixels in each column and use two threshold values to determine the values in a length-3 vector for each image. This vector represents the image as 1-1, 0, or $1$, based on the comparison with the thresholds.
A neural network with a hidden layer of depth 1, width 2, and the Sigmoid activation function $\sigma(x)=\frac{1}{1+e^{-x}}$ is used for classification. We approximate the sigmoid activation function with a piecewise constant function. The breakpoints in this experiment are $\{-8,-4,0,4,8\}$, and the value of the piecewise linear function in each interval is equal to the function value of the sigmoid function at the midpoint of that interval.
F.2 Robustness of hybrid approach: impact of coefficient matrix rounding
To investigate the convergence behavior of the hybrid algorithm in neural network training under different matrix rounding precision levels, we systematically rounded the matrix coefficients to digit precisions ranging from 1 to 10 decimal places. The algorithm was then executed for each precision setting until convergence to the optimal solution was achieved. Specifically, matrix elements were rounded to 1 through 10 significant digits.
Figure 6(a) displays the QUBO matrix evolution using circular heatmaps, highlighting the impact of precision levels on iterative trajectories. Despite the differences revealed in these trajectories, our results confirm that all trajectories converge to the optimal solution, consistent with the theoretical convergence of the QCGD algorithm. Notably, even under minimal precision conditions (e.g., 1-digit rounding), matrix adjustments in later iterations remain marginal. These consistently small perturbations across all precision levels suggest that a lazy evaluation strategy for QCGD could be viable. Such an approach would minimize redundant quantum circuit executions, thereby improving computational efficiency on quantum hardware.
The smoothed absolute differences of the objective value and constraint residual relative to the 10-digit precision baseline were recorded, with results visualized in Figure 6(b) and 6(c). This setup allowed us to assess the impact of numerical precision on convergence trajectories. Both subplots demonstrate that differences across all precision levels remain small, typically less than $10^{-2}$, indicating the hybrid algorithm’s robustness to precision variations. A downward trend is observed in the difference of constraint residual, with differences decreasing sharply after initial fluctuations and stabilizing post-1500 iterations, confirming convergence to a consistent optimal solution regardless of precision, as supported by prior results. The rapid initial decline in the constraint residual difference, particularly for lower precisions (e.g., 1-3 digits), suggests early sensitivity to numerical accuracy, while the differences remain stably small in later iterations. The near-overlap of higher precision curves (e.g., 7-9 digits) highlights diminishing returns from increased precision.
F.3 Analysis of Randomized Rounding Procedure
Let $\bar{x}\in[0,1]^{n}$ be the DLBO solution and denote by $F_{\mathrm{DLBO}}$ its objective value. Consider $K$ independent randomized roundings of $\bar{x}$,
$X^{(1)},\dots,X^{(K)},\qquad X^{(k)}_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j})\ \text{independently for }j=1,\dots,n,$
and write $F^{(k)}:=F(X^{(k)})$ as well as $F_{\min}:=\min_{1\leq k\leq K}F^{(k)}$.
As the objective $F$ is defined as the sum of per–sample losses measuring the discrepancy between the predicted labels and the true labels, we claim that $F$ is bounded and thus satisfies the bounded–difference property, i.e., there exist constants $\Delta_{1},\dots,\Delta_{n}$ such that changing only the $j$-th coordinate of the argument changes $F$ by at most $\Delta_{j}$. McDiarmid’s inequality [77] then yields, for any $t>0$ and each fixed $k$,
$\Pr\!\left(F^{(k)}\geq\mathbb{E}[F(X)]+t\right)\;\leq\;\exp\!\left(-\frac{2t^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right).$
Using the independence of the $K$ samples, we obtain
$\Pr\!\left(F_{\min}\geq\mathbb{E}[F(X)]+t\right)=\Pr\!\left(\bigcap_{k=1}^{K}\bigl\{F^{(k)}\geq\mathbb{E}[F(X)]+t\bigr\}\right)\leq\left[\exp\!\left(-\frac{2t^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right)\right]^{K}=\exp\!\left(-\frac{2Kt^{2}}{\sum_{j=1}^{n}\Delta_{j}^{2}}\right).$
Setting the right-hand side equal to $\delta\in(0,1)$ and solving for $t$ gives
$F_{\min}\;\leq\;\mathbb{E}[F(X)]+\sqrt{\frac{1}{2K}\Bigl(\sum_{j=1}^{n}\Delta_{j}^{2}\Bigr)\log\frac{1}{\delta}}\quad\text{with probability at least }1-\delta.$
We can assume that the discrete objective $F$ is a quadratic polynomial with zero diagonal, i.e.,
$F(x)\;=\;x^{\top}Qx+c^{\top}x,\qquad x\in\{0,1\}^{n},$
where $Q\in\mathbb{R}^{n\times n}$ is symmetric with $Q_{ii}=0$ for all $i$ and $c\in\mathbb{R}^{n}$. Because for binary variables $x_{i}^{2}=x_{i}$, so each diagonal coefficient $Q_{ii}$ can be absorbed into the linear term. Let $\bar{x}\in[0,1]^{n}$ be the relaxed solution and consider the independent Bernoulli rounding
$X_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j}),\quad j=1,\dots,n,$
with all coordinates independent. Then $\mathbb{E}[X_{j}]=\bar{x}_{j}$ and, for $i\neq j$,
$\mathbb{E}[X_{i}X_{j}]\;=\;\mathbb{E}[X_{i}]\mathbb{E}[X_{j}]\;=\;\bar{x}_{i}\bar{x}_{j}.$
Using these identities, we can compute
$\mathbb{E}[F(X)]=\mathbb{E}[X^{\top}QX]+\mathbb{E}[c^{\top}X]=\sum_{i,j}Q_{ij}\,\mathbb{E}[X_{i}X_{j}]+\sum_{j}c_{j}\,\mathbb{E}[X_{j}].$
Because $Q_{ii}=0$, the quadratic term reduces to
$\sum_{i\neq j}Q_{ij}\,\mathbb{E}[X_{i}X_{j}]=\sum_{i\neq j}Q_{ij}\,\bar{x}_{i}\bar{x}_{j}=\bar{x}^{\top}Q\bar{x},$
and the linear term is
$\sum_{j}c_{j}\,\mathbb{E}[X_{j}]=\sum_{j}c_{j}\bar{x}_{j}=c^{\top}\bar{x}.$
Therefore,
$\mathbb{E}[F(X)]=\bar{x}^{\top}Q\bar{x}+c^{\top}\bar{x}=F(\bar{x}).$
Substituting this bound into the previous display yields the comparison between the best-rounded solution and the DLBO objective value, i.e.,
$F_{\min}\;\leq\;F(\bar{x})+\sqrt{\frac{1}{2K}\Bigl(\sum_{j=1}^{n}\Delta_{j}^{2}\Bigr)\log\frac{1}{\delta}}$
with probability at least $1-\delta$.
Figure 7: Schematic of the setup of the coherent Ising machine (CIM).
Appendix G Principles of Coherent Ising Machine
As shown in Figure 7, we implement the CIM, an optical system engineered to address combinatorial optimization problems through optical parametric oscillators (OPOs). A continuous wave (CW) laser at 1560 nm initiates the process, feeding into a pump pulse preparation stage. Here, intensity modulators (IM1, IM2, IM3), erbium-doped fiber amplifiers (EDFA1, EDFA2), and filters shape and amplify the light into pump pulses. These pulses enter a second harmonic generation (SHG) unit, producing a frequency-doubled signal, which is then directed into a periodically poled lithium niobate (PPLN) crystal. The PPLN facilitates phase-sensitive amplification, forming a degenerate OPO that generates coherent signal and idler photons. These photons are measured via balanced homodyne detection (BHD), with an FPGA module providing feedback to steer the system toward optimal solutions by mapping the problem onto the Ising model’s energy landscape.
CIM operates by simulating the Ising Hamiltonian, defined as
$H=-\sum_{i,j}J_{ij}\sigma_{i}\sigma_{j}-\sum_{i}h_{i}\sigma_{i},$ (139)
where $\sigma_{i}\in\{1,-1\}$ represents spins, $J_{ij}$ denotes coupling strengths, and $h_{i}$ accounts for external fields. In the PPLN crystal, pump pulses drive parametric amplification, creating optical pulses that encode the spins as two possible phase states. The BHD measures the pulses’ phase and amplitude, determining the system’s energy state. The FPGA uses this data to adjust parameters like pump intensity, effectively minimizing the Hamiltonian through feedback. A chopper and intensity modulators ensure stable pulse circulation in the loop, while the feedback introduces coupling between pulses, mimicking the $J_{ij}\sigma_{i}\sigma_{j}$ interaction terms.
The system’s dynamics allow it to explore the Ising energy landscape efficiently. The optical pulses evolve in parallel, with their interactions governed by the feedback loop, driving the system toward the ground state of the Hamiltonian, which corresponds to the optimization problem’s solution. The amplifiers and filters maintain pulse quality, while the PPLN’s nonlinear efficiency ensures robust signal generation. The BHD’s noise suppression enhances measurement precision, enabling the CIM to handle complex problems by leveraging the coherence and parallelism of optical signals.